A Bibliography of Publications about the SPSS (Statistical Package for the Social Sciences) System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 June 2023 Version 1.36

Author/editor word crossreference

Abd [424]. Abouserie [380]. Academic [99]. Ahlgren [143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 177, 178]. Akron [98]. al-Razzaq [424]. Alan [332, 441, 212]. Allen [162, 524]. American [24]. Amsden [80, 102]. Andrew [143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 177, 178, 500]. Angie [503]. Anonymous [26, 163, 164, 165, 179, 180, 181, 182, 183, 184, 185, 186, 209, 210, 272, 306, 307, 308, 309, 331, 406, 504, 506, 13, 25, 129, 130, 36, 37]. Anthony [219]. Arigo [187]. Arizona [75].

Arkkelin [344]. Association [24]. Austin [53].

Babbie [407, 440]. Backhaus [310]. Badarinathi [408]. Barbara [6]. Barcikowski [166]. Barry [312, 313, 43]. Basto [508]. Bauer [188]. Beadle [77, 88]. Becker [227]. Belle [79]. Ben-Liou [38]. Benjamin [207]. Benson [78, 189]. Bent [1, 2, 3]. Berenson [169]. Berkeley [133, 142]. Bernard [39]. Bertha [190]. Betty [78, 189]. Beutel [100]. Blake [364]. Bloomberg [79]. Bob [11]. Bodendorf [121]. Brodeur [381, 409]. Brosveet [27]. Bruce [16, 123]. Bryman [332, 441]. Butler [190].

California [133, 142]. Cathy [193]. Center [99, 53, 103, 51, 75, 131, 15, 98, 204, 205, 499].

Charles [43]. Charlick [14]. Chongzi [192]. Christian [384]. Christiansen [101]. Chung [86]. City [51]. Cleary [80, 102]. Clements [54]. Clevie [128]. Clifford [227]. Clogg [227]. Colin [385, 444]. Computation [53]. Computer [99, 103, 90, 98, 51, 75, 204, 205]. Computing [32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117, 35, 23, 131, 208, 133, 142]. ComTrain [273, 255]. Concepcion [10]. Corporation [255]. Cramer [332, 441, 442]. Crisler [344].

Dag [274, 311]. Dale [1, 2, 3]. Dallas [514]. Dan [34, 47]. Daniel [344, 52]. David [524, 79, 502, 11, 169, 344]. DeBoer [167]. Del [226]. Dennis [49, 50]. Diane [220]. Dipl.-Sozialwirt [188]. Doerr [168]. Dometrius [382]. Don [11]. Donald [167]. Douglas [167]. Downing [16]. Drake [207]. Drone [81]. Duncan [332, 441, 442].

Earl [407, 440]. Ebeling [82]. Edward [123]. Edwards [52]. Edwin [511]. Eelko [415]. Einstein [256, 228]. Elaine [256, 228]. Elizabeth [103]. Ellen [54]. Evers [167].

Fan [443]. Favero [226]. Federer [139]. Felix [188]. Fetcher [103]. Field [41]. Fischer [410]. Fisher [414]. Foster [383, 411]. Fred [440]. Frederick [167]. Freimut [121]. Fridlund [212]. Friesen [364]. Frude [257, 258, 412, 413].

Garfield [167]. Gary [105]. Geert [511]. George [43]. Georgia [32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117]. Gerhard [276]. Gilles [256, 228, 48]. Gladney [128]. Glenn [417]. Godfrey [509]. Graham [107]. Gray [385, 444]. Green [5]. Greenberg [123]. Gregory [507]. Gustav [366]. Gutwein [104].

Hadi [236]. Hadlai [77, 84, 124, 125, 17, 1, 2, 3, 4, 44, 45, 46, 58, 59, 87, 107, 137]. Halley [440]. Harry [314, 135, 138, 195]. Hays [10, 55, 105]. Hedderson [259, 365, 414]. Helen [168]. Heller [143, 144, 145, 146, 147, 148, 149, 150, 151, 158, 159, 160, 161, 177, 178]. Helmut [254]. Helmuth [215]. Henry [410, 214, 6]. Herbert [43]. Heuvel [511]. Hinkley [103]. Ho-jong [416]. Hohlen [78, 189, 83, 84]. Holm [458]. Holzer [384]. Hossain [517, 518]. Howe [187]. Hrin [28]. Hubert [85, 127]. Hudson [139]. Huizingh [415]. Hull [77, 84, 124, 125, 17, 1, 2, 3, 4, 44, 45, 46, 58, 59, 87, 107, 137]. Humphrey [126].

III [410]. Inc. [273, 255, 445, 446, 447, 90, 73, 91, 92, 140, 174, 175, 176, 196, 197, 198, 199, 200, 201, 202, 221, 222, 223, 224, 225, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 263, 264, 265, 266, 267, 268, 269, 270, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 345, 346, 347, 348, 349, 350, 351, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 376, 377, 378, 395, 396, 397, 398, 399, 400, 401]. Inc. [402, 403, 404, 405, 425, 426, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439, 450, 451, 452, 453, 454, 455, 460, 461, 462, 463, 464, 465, 466, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476, 477, 478, 480, 481, 482, 483, 484, 485, 486, 487, 488, 489, 490, 491, 492, 493, 494, 495, 496, 497, 498, 35]. Information [15]. Ingwersen [106]. Institute [32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117, 41]. Ivy [333].

Jakes [11]. James [106]. Jan [97]. Jane [170]. Jarle [27]. Jay [315, 387]. Jendrek [213]. Jeremy [383, 411]. Jianjun [514]. Jie [515]. Joaquim [505]. John [101, 104, 259, 365, 414, 226]. Johnson [514]. Jon [82]. Jonah [105]. Jonan [55]. Jonathan [509]. Jones [105]. Joreskog [274, 311]. Jorge [22, 68]. Jørgen [458]. José [508, 516]. Joseph [134, 502]. Joy [132]. Jr. [74]. Justin [5].

Kahler [275]. Kansas [99]. Karpman [190]. Kathy [52]. Keith [42]. Kellough [214]. Ken [305]. Kershner [71, 119]. Keyes [312, 313]. Kilman [394]. Kim [416]. Kimble [107]. King [82]. Kinnear [385, 444]. Klaus [310]. Klecka [17]. Kockläuner [276]. Kolles [314]. Kosfeld [386]. Kuffner [215]. Kwok-Leung [305].

Laboratory [23]. Laforge [85, 127]. Lamp [134]. Larry [344]. Larsen [333]. Larson-Albers [193]. Lee [86, 333]. Lei [192]. Lemmon [502]. Leuthold [11]. Levine [169, 366]. Lewis [123]. Library [24]. Lincoln [103]. Linda [128]. Ling [135, 138]. Link [167]. Linn [456]. Longford [457]. Loots [509]. Lynch [500]. Lynn [220].

Macbriar [29]. Magidson [315, 387]. Manfred [384, 141]. Manuel [508, 516]. Marchant [42]. Margaret [213]. Marija [89, 137, 171, 172, 173, 194, 216, 217, 218, 229, 230, 231, 232, 233, 234, 235, 260, 261, 262, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 316, 334, 335, 336, 337, 338, 339, 340, 341, 342, 343, 367, 368, 369, 370, 371, 372, 373, 374, 388, 389, 390, 391, 392, 393, 418, 419, 420, 421, 422, 423, 445, 446, 447, 448, 449]. Mário [508]. Mark [227, 55, 105, 169]. Marques [505]. Martine [28]. Maurice [42]. Max [123]. Mayers [207]. Maykovich [333]. McNeil [11]. McNett [43]. Md [517, 518]. Media [103]. Melanie [132].

Melinda [414]. Meyer [417]. Michael [80, 102, 82]. Milan [28]. Millikan [170]. Mills [55, 105]. Mina [78, 189, 83, 84]. Minako [333]. Minnesota [103, 204, 205]. Minnesota.Computer [499]. Mitchell [190]. Miyata [52]. Molenberghs [511]. Morrison [136]. Mosier [143, 144, 145, 146, 147, 148, 149, 150, 151, 158, 159, 160, 161, 177, 178]. Moyazzem [517, 518]. Murdock [43].

N. [22, 68]. Nancy

[143, 144, 145, 146, 147, 148, 149, 150, 151, 158, 159, 160, 161, 177, 178, 136]. Nathan [42]. Nazanin [511]. Neil [257, 258, 412, 413, 512, 271]. Nelson [382]. Newcastle [23]. Nick [457]. Nie [77, 14, 84, 124, 125, 17, 1, 2, 3, 4, 8, 9, 18, 19, 20, 21, 30, 31, 44, 45, 46, 56, 57, 58, 59, 60, 87, 88, 107, 108, 137]. Nielsen [501]. Nocks [256, 228]. Nooraee [511]. Norman [77, 14, 84, 124, 125, 17, 1, 2, 3, 4, 8, 9, 18, 19, 20, 21, 30, 31, 44, 45, 46, 56, 57, 58, 59, 60, 87, 88, 107, 108, 137]. Normand [409]. Northwestern [131]. Norusis [89, 137, 171, 172, 173, 194, 216, 217, 218, 229, 230, 231, 232, 233, 234, 235, 260, 261, 262, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 316, 334, 335, 336, 337, 338, 339, 340, 341, 342, 343, 367, 368, 369, 370, 371, 372, 373, 374, 388, 389, 390, 391, 392, 393, 418, 419, 420, 421, 422, 423, 445, 446, 447, 448, 449].

Office [32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117]. Ortiz-Marcos [516].

Padua [22, 68]. Page [6]. Patrice

[381, 409]. Patricia

[143, 144, 145, 146, 147, 148, 149, 150, 151, 158, 159, 160, 161, 177, 178]. Patrick [214, 207]. Paul [79, 81, 10, 385, 444, 71, 49, 50, 93, 94, 95, 96]. Pereira [508]. Peter [100, 43, 226].

Petersen [458]. Platt [213]. Pooler [69]. Precht [384, 34, 47]. priemyselná [28]. Prime [90]. Privitera [507]. Processing [15].

Ramalingam [513]. Ranger [409]. Rauch [49, 50]. Ravija [408]. Reda [380]. Reinhard [215]. Reinhold [386]. Research [24]. Resources [103]. Reynolds [6]. Rhonda [52]. Richard [333]. Robert [166, 14, 80, 102, 510, 135, 138, 379]. Roberts [135, 138, 195]. Ronald [71]. Round [24]. Rowland [344].

Sá [505]. Sabroe [459]. Salavitabar [236]. Salkind [512]. Sam [81]. Schafer [502]. Schubo [375]. Searle [71, 119, 139]. Services [133, 32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117, 142, 208]. Shanmugam [513]. Shayle [71, 119, 139]. Shimizu [10]. Shin [394]. Shu-yüan [206]. Shurbaji [424]. skola [28]. Smith [42]. Sola [219]. Sorbom [274, 311]. Spaulding [220]. SPSS [445, 446, 447, 73, 91, 92, 140, 174, 175, 176, 196, 197, 198, 199, 200, 201, 202, 221, 222, 223, 224, 225, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 263, 264, 265, 266, 267, 268, 269, 270, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 345, 346, 347, 348, 349, 350, 351, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 376, 377, 378, 395, 396, 397, 398, 399, 400, 401, 402, 403, 404]. SPSS [405, 425, 426, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439, 450, 451, 452, 453, 454, 455, 460, 461, 462, 463, 464, 465, 466, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476, 477, 478, 480, 481, 482, 483, 484, 485, 486, 487, 488, 489, 490, 491, 492, 493, 494, 495, 496, 497, 498]. State [208]. Steele [134]. Stene [226]. Stephan [169]. Stephen [134]. Steven [219]. Stirling [42]. Stredná [28]. strojnícka [28]. Susan [69]. Svend [459]. Systems [35].

Table [24]. Technology

[32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117]. Tellier [48]. Texas [53]. Thaller [141]. Theodor [509]. Theodore [139]. Thistlethwaite [79, 81, 49, 50, 93, 94, 95, 96]. Thomas [69, 76]. Timothy [10]. Tine [501]. Tompkins [128]. Trollvik [97]. Trotter [74]. Tyne [23].

United [35]. University

[99, 53, 133, 103, 142, 51, 75, 23, 208, 98, 51, 75, 204, 205, 499, 131]. upon [23].

Vasco [255]. ViAnn [77, 88]. Vogelback [131].

Wade [503]. Walter [139]. Wang [514]. Webb [187]. Weber [208]. Weiss [271]. Weng [206]. Werner [375]. West [379]. Wilke [254]. William [162, 123, 17, 69, 74]. Wilson [10]. Wittenberg [215]. Wolf [76, 132]. Wolf-Michael [275]. Wong [305]. Wright [207].

Yao [38]. Yihua [443]. York [51].

Zaharris-Esgro [52]. Zhao [515]. Zimmerly [79].

Title word cross-reference

$104.00 [518]. $130.00 [518]. $44.95 [503]. T M [184]. ^{X} [165, 224, 256].

0 [457, 500, 501, 458, 459]. 0-340-76112-1 [501]. 0-387-94821-X [457]. 0-387-98563-8 [500]. 0-471-33869-9 [458]. 0-7169-6927-0 [500]. 0-7169-7385-0 [500]. 0-8039-9055-3 [459].

1 [501, 503]. 1-2-3 [364]. 1-58488-369-3 [503]. 10-14 [522]. 11 [183, 136, 107, 92]. 11-44 [226]. 1985 [524].

2 [518, 281, 286, 339, 341, 343, 242, 287, 288, 289, 295, 318, 329, 345, 354, 355, 356, 359, 360, 484]. 24.80 [406]. 2nd [501].

3 [459, 513, 503]. 3-8252-1602-0 [406]. 3-89028-132-X [386].

5.0 [415].

6.0/CDC6000 [18]. 6.1 [446, 452, 455].

7-9 [125]. ’78 [520].

8 [100]. 8-10 [500]. ’81 [521]. ’82 [522]. 8A [73].

9 [458]. 95 [502]. 978 [518, 513]. 978-1-119-58304-2 [518]. 978-1-4522-1648-3 [513].

A. [506]. Acceptance [515]. Accessing [193, 196, 251, 236]. Accompany [14, 271]. ACO [139]. acquisition [220]. Adding [224, 223]. Additional [260]. Advanced [216, 217, 229, 233, 234, 281, 334, 335, 340, 388, 391, 418, 421, 445, 325, 425, 434, 467, 468, 469, 487, 339, 293]. Adventures [407, 440, 459]. Advertising [515, 79]. Advisor [314]. agencies [132]. Agriculture [128]. aid [76]. al [424]. al-Arabiyah [424]. al-duwal [424]. al-Iqtisad [424]. al-qiyasi [424]. al-tatbiqi [424]. ala [424]. algorithm [456]. Algorithms [89, 172, 73, 222, 248, 378]. Alistair [500]. Allen [394]. Analyse [85, 127, 254, 39]. Analyses [504, 71, 503]. Analysis [380, 407, 440, 508, 332, 441, 102, 167, 383, 411, 275, 276, 215, 366, 135, 500, 230, 231, 260, 277, 278, 337, 367, 516, 458, 138, 344, 428, 429, 459, 379, 254, 144, 146, 148, 151, 152, 153, 154, 155, 156, 157, 161, 177, 178, 163, 192, 510, 106, 312, 313, 457, 69]. analysis/graphics [313]. Andel [28]. Anfanger [275]. Angewandte [276]. Anguish [458]. Annotated [71, 139]. Annotations [166, 177]. Announces [272, 163, 181, 306, 331]. Annual [519, 520, 521, 522]. ANOVA [384, 109, 71, 119]. Answers [14]. Anwender [121]. applicable [187]. Application [26, 81, 82, 79, 312, 94, 95, 132]. Applications [213, 274, 311, 192, 517, 518, 127, 51]. Applied [517, 505, 394, 276, 518]. Approach [81, 417, 79, 95, 96]. April [56]. Arabiyah [424]. Arbeitsmethode [141]. archives [82]. Arnold [501]. Articles [214]. assisted [132]. Atlas [43]. Auswertung [314]. Auswertungen [314, 215]. Auto [162]. Auto-Mode [162]. available [180, 184, 185, 82, 123].

B [457, 500]. B6700 [7]. Babbie [459]. BALANOVA [167]. Base [286, 336, 342, 389, 392, 419, 422, 446, 358, 397, 398, 403, 426, 435, 455, 470, 471, 488, 443, 341, 448, 298, 402]. Based [344, 515]. Basic [380, 101, 169, 137, 103]. Basics [386, 201]. Batch [169, 44, 45, 46, 59, 107, 228, 93, 136, 57, 140, 38]. Beginner [383, 411, 500]. Beginners [275]. Behavioral [344, 507]. Beispiele [386]. Beispielen [314]. Beschreibung [100]. Between [514, 18]. bi [424]. bi-al-istikhdam [424]. Bibliography [214]. Binary [516]. biologiques [85, 127]. bivariate [517, 518]. Blind [509]. BMD [85, 127]. BMDP [166, 39, 167, 384, 215, 85, 127]. Boca [503]. Book [406, 504, 506, 518, 386, 457, 500, 501, 458, 459, 513, 503, 41]. box [309, 243, 266]. Brian [503]. Brosius [386]. Building [513]. bull [212]. Bulletin [83]. bundling [309]. Burroughs [7, 52]. Business [169, 369, 373, 394].

ca [386]. Cal [211]. calculate [456]. Calculations [442]. calibration [187]. California [521]. capabilities [443]. Capture [251, 302]. Carolina [10]. case [456]. Categories [348, 472, 473]. CDC [53, 133, 49, 50, 142]. CDC6000 [18]. Census [174, 200]. Center [103, 106]. CHAID [315, 387, 427, 474, 491]. Chapman [503]. charyo [416]. Chicago [520, 523]. Ciencias [22, 68]. Cleaning [143]. Climatic [187]. CMS [224, 351, 478, 129, 203]. Coakes [458]. Code [224, 41, 223]. Codebook [377, 475]. Coding [10]. Collins [106]. Command [92]. commands [103]. communications [209]. Compact [375]. Company [386]. Comparaison [39]. Comparative [167]. Comparing [511]. Comparison [123, 443, 384]. COMPUSTAT [236]. Computer [524, 166, 211, 442, 106, 213, 314, 11, 344, 187, 294, 296, 297, 325, 326, 71, 139, 220, 97, 131, 76, 132, 524]. Computer-Based [344]. Computer-Supported [314]. computergestützte [314]. Computers [118, 185, 56]. Computing [379]. Concepts [380]. Concise [364]. condiciones [22, 68]. Conference [519, 520, 521, 522, 523]. consulting [181]. contingency [147]. control [86, 98]. Conversational [58, 87, 108, 138, 207]. Conversion [29, 133, 142]. Coordinator [320]. Coordinators [519, 520, 521, 522]. COPY [460]. Corporation [92]. Course [23]. COV [139]. covariance [144, 146, 139]. Cramer [406]. Cray [272]. CRC [503]. CSUC [82]. Cyber [97]. Cyberbullying d [100]. Daniel [518]. Data [380, 504, 407, 440, 332, 441, 102, 383, 411, 10, 415, 275, 135, 500, 29, 230, 231, 260, 277, 278, 337, 367, 69, 138, 344, 174, 237, 245, 246, 263, 287, 290, 318, 319, 396, 428, 429, 431, 461, 464, 466, 476, 459, 141, 143, 151, 152, 153, 154, 155, 156, 157, 177, 178, 179, 187, 82, 443, 41, 123, 510, 313, 511, 197, 200, 291, 236, 71, 103, 215, 457, 513]. databases [307]. DATATRIEVE [193, 196, 251, 302]. Datenanalyse [188, 406]. Datenanalysessyteme [215]. Datenanalysesystem [275]. Datenverarbeitung [141]. Dbase [424]. DBMS [460]. DBMS/COPY [460]. Debugging [74]. DEC [180, 136, 107, 198]. December [59]. decision [312]. DECsystem [25]. DECsystem-10 [25]. Delaware [6]. Demo [265]. Denis [518]. Department [128]. Description [100]. Design [166]. Developer [437]. Developing [502]. Development [128, 132, 305, 310]. dictionary [179]. Differences [18]. Digital [92, 223, 193, 196, 140, 291]. directed [103]. Discrepancies [514]. diskette [459]. diskettes [458]. distributed [384]. Division [128]. DM [406, 386]. Doctors [314]. Documentation [70]. documenting [76]. documents [19]. Doing [134, 500]. Doktoranden [314]. d’ordinateur [85]. Dryden [408]. d’utilisation [448]. duwal [424]. Dvadsatpät [28].

e-book [518]. each [456]. Economic [128]. Economics [394]. EDIT [170]. edited [504]. Edition [500, 271]. edn [501]. Education [207]. Educational [380, 516]. effectiveness [220]. efficient [74]. Einführung [275, 314, 386, 215, 141]. Elmer [170]. Entering [246, 287, 461]. Entry [415, 245, 263, 287, 290, 318, 319, 396, 461, 464, 466, 476]. Entwicklung [310]. Environments [516]. Equipment [92].

Estadístico [22, 68]. Ethnographic [43]. Evaluation [314]. Evaluations [314, 215]. Everitt [503]. Examination [514]. Examples [167, 314]. Excel [424]. Exercises [14]. Explanation [10]. Explanations [36]. exploration [192]. EXPORT [179]. Extended [92].

F [459]. F17703 [41]. Facilities [77, 125, 88, 175]. Facility [321, 179]. Factor [508, 148]. Factorial [144]. Fall [26]. Fast [273, 255, 295]. features [212]. Featuring [244, 327]. Feed [515]. Fetcher [103]. fichiers [48]. fichiers-SPSS [48]. Fifth [521]. File [81, 43, 237, 79, 41, 197, 291, 294, 296, 297, 325, 326, 95, 96]. files [179, 103]. Finance [81]. first [163]. Fischer [406]. flexible [78]. Forecasting [135, 195]. Forge [459]. Formats [237, 197, 291]. Fort [106]. Fortran [502]. Forum [24]. Foster [500]. fot [28]. FOUBAR [67]. four [212]. Francisco [521].

G [386, 458, 10]. G3SLS [110]. GCOS [99, 184]. GEE [511]. geepack [511]. generation [38]. generator [78]. GENLIN [511]. GENMOD [511]. GENSTAT [39]. Gerber [457, 500]. Getting [288, 289, 317, 318, 345, 395, 462, 463, 464, 465]. going [133, 142]. graf [28]. Grancic [28]. graph [309, 243, 266]. graph-in-the-box [309, 243, 266]. graphic [386]. Graphics [221, 238, 239, 241, 242, 244, 320, 321, 322, 327, 349, 350, 356, 376, 399, 400, 405, 477, 498, 312, 313, 249, 401, 307, 498]. graphs [513]. group [181]. Guide [380, 14, 54, 383, 411, 364, 412, 413, 124, 274, 311, 333, 366, 135, 500, 30, 108, 137, 171, 173, 216, 230, 231, 260, 277, 278, 279, 334, 335, 336, 337, 338, 367, 389, 392, 419, 422, 446, 70, 90, 12, 91, 176, 252, 301, 346, 351, 352, 353, 354, 355, 361, 362, 363, 397, 398, 403, 426, 428, 429, 431, 435, 452, 455, 467, 469, 470, 471, 478, 480, 481, 483, 484, 488, 490, 495, 496, 497, 375, 394, 37, 187, 122, 191, 133, 257, 258, 5, 105, 457, 170, 33, 49, 50, 142, 52, 501]. Guided [265]. Gustav [406].

H [100]. Håndbok [97]. Hall [500, 503]. Hall/CRC [503]. Halley [459]. Hamburg [386]. Handbook [162, 102, 510, 11, 503, 504]. Hans [406]. hardcover [518]. Harvard [244, 327, 405, 498]. Hassett [271]. hate [512]. health [132]. Help [404, 492]. Helper [80]. Hewlett [186]. Hill [386]. Historians [141]. Historiker [141]. Hoboken [518]. Honeywell [99, 184, 56]. horns [212]. Hotel [524]. Howard [500]. Hrin [28]. hwaryonghan [416]. Hybrid [515].

i.e [28]. IBM

[182, 133, 410, 510, 194, 218, 229, 232, 234, 235, 281, 282, 286, 339, 341, 343, 197, 224, 237, 241, 242, 245, 246, 250, 253, 264, 267, 269, 287, 288, 289, 290, 293, 295, 298, 318, 328, 329, 345, 351, 352, 356, 357, 359, 360, 478, 480, 142]. IBM-PC [410]. IBM/MVS [480]. IDA [135, 138, 195, 207]. Idea [310]. IFT [381, 409]. II [263, 287, 290, 299, 318, 396, 461, 464, 466, 476]. III [424]. Illinois [523]. Illustrated [501]. Implementation [83]. IMPORT [179]. IMPORT/EXPORT [179]. Imputation [514]. Inc. [163, 164, 181, 307, 219]. increases [309]. Information [515, 163]. informatique [381, 409]. Informix [306, 307]. injury [456]. Input [166]. Installing [396, 466]. Instructional [26]. Instructions [260]. Instructor [231, 371]. instruments [183]. Interactive [79, 81, 135, 94, 95, 96, 49, 50, 93, 76, 38]. Intercultural [516]. Interface [524, 305, 76]. interfaces [306, 307]. interpretacion [68, 22]. Interpretation [504]. Introduces [308, 164, 307]. Introducing [442]. Introduction [145, 408, 364, 415, 275, 314, 215, 138, 226, 141, 499, 16]. Introductory [11, 137, 171, 173, 279, 338, 12, 481, 495, 271]. Iowa [5]. Iqtisad [424]. iqtisadiyat [424]. ISBN [406, 518, 386, 457, 500, 501, 458, 459, 513, 503]. ISBN-13 [513]. ISS [456]. ISSUE [520, 521, 522, 523]. istikhdam [424]. IV [24].

J [518, 500, 458]. Ján [28]. JCL [86]. Jeremy [500]. JFACTOR [61]. Job [98, 103, 86]. John [518]. July [523].

Kentucky [524]. Kermit [209]. Kerr [500]. Keywords [120]. Kit [437]. kol [28]. Kompakt [375]. Konzeption [310]. Kozub [500]. Kristin [500]. kwahak [416].

L [458]. Lab [211]. Landau [503]. language [86, 98]. Learning [256, 228]. level [99, 5, 56]. Lexington [524]. Libraries [10]. Library [42]. Lincoln [103]. Linear [227]. LISREL [163, 274, 311, 202, 304, 346, 482, 430]. list [40, 53]. Log [227]. Log-Linear [227]. logiciels [39, 127]. Logistic [516]. London [500, 501]. longitudinal [511]. Lotus [364]. Louisiana [522].

Machine [305]. Macintosh [323, 331, 353, 483]. Made [259, 365, 414, 385, 444]. mainframe [209]. Man [305]. Man-Machine [305]. management [443, 123, 69]. management/manipulation [443]. manipulation [443, 103]. MANOVA [111]. Manual [410, 1, 2, 4, 8, 9, 31, 44, 45, 46, 58, 59, 87, 231, 286, 371, 34, 47, 358, 7, 35, 205, 13, 189, 14, 101, 84, 18, 57, 341, 298, 330]. Manuals [214, 19]. manuel [448]. many [134]. Map [347]. Mapping [264]. March [524, 57]. Marketing [310]. Marketing-Konzeption [310]. Martin [28]. Master [43]. Mathematical [51]. MATLAB [505]. matricielle [85]. Maze [213]. McGraw [386]. McGraw-Hill [386]. med [27]. Medical [314]. Medicine [314]. Medizin [314]. Menu [508]. Method [515]. Methodenbankrahmen [121]. Methods [380, 141, 220]. Micro [209]. microform [76, 305]. Milan [28]. Mini [14]. Mini-manual [14]. Minimalist [417]. Minitab [509, 384, 379, 364]. Mode [162]. Modeling [227]. models [384]. MTS [126, 34, 47]. MTS/SPSS [126, 34, 47]. Muenchen [506]. MULT [112]. multgee [511]. Multidimensional [147]. Multiple [150, 384, 514, 149]. Multivariate [146, 504, 192, 510, 517, 518]. multivariée [85, 127]. MUSIC [122, 191]. MVS [224, 352, 480].

Nach [100]. natural [517, 518]. navrhol [28]. Newsletter [72, 479]. Nie [100, 100]. NJ [518]. No [457, 500, 501, 458, 503, 41]. NONLINEAR [62]. Nonuser [10]. normally [384]. Norman [100, 100]. North [10]. Notebook [320]. Notes [99, 56, 381, 409]. NPAR [63]. NUCC [19]. numbers [212]. Numerical [141]. Numerische [141]. Numerous [314]. Nutzung [254]. NWA [364].

ob [28]. October [520, 521, 522]. ODC [40]. oferta [22, 68]. one [161]. one-way [161]. Online [162, 26, 55, 35, 159, 105, 40, 64, 204]. Operations [137, 351, 352, 353, 354, 355, 362, 363, 478, 480, 483, 484, 490]. option [309]. Ordinal [508, 511]. oriented [38]. Orleans [522]. OS/2 [354, 355, 484]. Outline [121]. Output [166, 71, 139].

paced [93]. Package [13, 187, 121, 14, 16, 82, 103, 5, 104, 214, 17, 11, 1, 2, 3, 4, 8, 20, 21, 30, 60, 69, 70, 12, 6, 92, 431, 141, 49, 50, 98, 7, 35, 37, 331, 76]. Packages [166, 364, 254, 212, 123]. Packard [186]. pages [457, 500, 501, 458, 459, 503].

Paguete [22]. pamätnica [28]. Paquete [68]. para [22, 68]. Part [446, 455, 499]. partie [409]. Pasant [501]. Payment [330]. PC [182, 380, 182, 209, 308, 309, 310, 381, 409, 211, 273, 255, 410, 383, 411, 364, 257, 258, 412, 413, 365, 415, 385, 333, 456, 315, 387, 194, 217, 218, 229, 232, 233, 234, 235, 260, 261, 262, 278, 280, 281, 282, 283, 284, 285, 286, 316, 339, 340, 341, 342, 343, 367, 368, 371, 372, 373, 374, 389, 390, 199, 240, 241, 242, 243, 244, 245, 246, 250, 253, 264, 265, 266, 267, 268, 269, 287, 288, 289, 290, 292, 293, 294, 295, 296, 297, 298, 318, 324, 325, 326, 327, 328, 329, 345, 347, 356, 357, 358, 359]. PC [360, 377, 395, 396, 397, 399, 400, 401, 402, 404, 405, 429, 431, 432, 433, 450, 462, 463, 466, 468, 474, 475, 482, 485, 486, 492, 498, 424, 219, 386]. PC/XT [194]. PC/XT/AT [229, 234, 281, 286, 339, 341, 343, 242, 253, 287, 288, 289, 293, 298, 318, 329, 345, 359, 218, 232, 235, 282, 328, 357, 241, 245, 246, 250, 264, 267, 269, 290, 295, 356, 360]. PCb [369, 370, 394]. PCp [416, 388, 247]. PDP [136, 107, 92, 226]. PDP-11 [136, 107, 92]. PDR3173 [70, 90]. people [512]. Perkin [170]. Perkin-Elmer [170]. Pine [459]. Planners [104]. Plaza [524]. PLOT [113]. Plus [368, 369, 370, 371, 372, 374, 460, 424, 358]. Pocket [54, 124, 91, 40]. POISE [226]. pollen [187]. Poly [211]. pour [39, 381, 409, 127, 448]. power [212]. pp [406, 518, 386]. Practical [151, 152, 153, 154, 155, 156, 157, 177, 178, 314, 141]. Practice [501]. Practitioners [509]. praktische [386]. praktischer [314]. praxisorientierte [141]. Preliminary [58, 87, 70, 73]. PRELIS [346, 482, 430]. présentation [85]. Press [459, 408]. Price [457, 458, 459, 503]. priemyselnej [28]. Prime [70, 90, 185, 118]. Primer [17, 170, 270]. Pro [424, 198]. Procedure [88, 163]. Procedures [77, 125, 175, 36, 514, 384]. Proceedings [524, 519, 520, 521, 522, 523]. Processing [174, 141, 200]. production [103]. Professional [390, 393, 420, 423, 447, 449, 436, 451, 485, 489, 198]. Program [100, 121, 10, 274, 311, 314, 29, 254, 38]. Programm [100, 60]. programmed [16]. Programmer [90]. programmes [85]. Programming [100, 30, 37]. Programmpaket [121]. Programmpakete [254]. Programmversionen [100]. Programs [224, 187, 443, 223, 131]. Pro^{T M} [164]. provide [212]. Provisional [1, 2]. PS [281, 286, 339, 341, 343, 242, 287, 288, 289, 295, 318, 329, 345, 356, 359, 360]. PS/2 [281, 286, 339, 341, 343, 242, 287, 288, 289, 295, 318, 329, 345, 356, 359, 360]. Psychological [380]. Psychologists [379]. Psychology [207]. Public [10]. Publication [73]. Publications [500]. Publishers [501]. punsok [416]. Puri [501].

QCA [515]. qiyasi [424]. Quantitative [332, 441, 141]. Quatro [424]. Quatro-Pro [424]. quick [75].

R [506, 509, 508, 517, 505, 511, 518, 506]. R-geepack [511]. R-Menu [508]. R-multgee [511]. R-repolr [511]. R. [504]. Radisson [524]. range [212]. Ratgeber [314]. Raton [503]. recording [24]. recreation [94]. Reference [34, 47, 346, 361, 398, 426, 452, 470, 40, 53, 75]. références [381, 409]. Region [10]. Regression [276, 138, 149, 150, 114]. Regressions [516]. Regressionsanalyse [276]. Reinhard [406]. relating [19]. Release [99, 54, 83, 124, 31, 44, 45, 46, 58, 59, 87, 89, 171, 279, 337, 391, 392, 393, 418, 419, 420, 421, 423, 70, 73, 91, 175, 239, 322, 349, 350, 355, 398, 403, 426, 428, 438, 439, 470, 471, 484, 495, 180, 84, 56, 57, 140]. Releases [77, 125, 88, 493]. RELIABILITY [33, 115]. repolr [511]. report [78, 116]. Reporting [158]. Rerun [321]. Research [380, 407, 440, 166, 442, 214, 42, 459, 254, 123]. RESPONSE [112]. resultados [22, 68]. results [158]. retailing [95]. Review [406, 504, 506, 509, 518, 386, 457, 501, 458, 459, 513, 219, 503, 212, 312]. reviewed [513]. Reviews [500]. rokov [28]. rul [416]. run [25].

S [457, 458, 503, 416, 369, 370, 388, 247, 394, 402]. S/PC [402]. Sabine [503]. Sage [500]. sahoe [416]. San [521]. SAS [506, 509, 166, 443, 364, 123, 384, 215, 127, 511]. SAS-GENMOD [511]. SAS^{T M} [514]. science [524]. Sciences [524, 100, 82, 4, 60, 69, 6, 344, 141, 49, 50, 98, 35, 381, 409, 85, 127, 507, 13, 14, 16, 103, 5, 104, 214, 17, 1, 2, 3, 8, 20, 21, 30, 70, 12, 92, 7, 37]. Scientists [187, 332, 441, 314, 517, 518]. score [456]. SCSS [58, 87]. Second [520, 271]. Selected [214]. Self [14, 23, 93, 76]. self-documenting [76]. self-paced [93]. Self-Teaching [14, 23]. SEM [515]. September [92]. Series [195, 99, 56, 131]. Service [330, 132]. Seventeenth [524]. severity [456]. Shaping [160]. shih [206]. shou [206]. Silche [416]. Simple [259, 365, 414, 385, 444, 103]. Singapore [458]. SIR [123]. Sixth [522]. skills [101, 220]. skoly [28]. Social [13, 187, 407, 440, 100, 332, 441, 14, 442, 382, 16, 82, 103, 410, 5, 104, 214, 17, 1, 2, 3, 4, 8, 20, 21, 30, 60, 69, 70, 12, 6, 344, 92, 459, 141, 49, 50, 98, 7, 35, 37, 254, 517, 518, 330]. sociales [381, 409, 85, 127, 68, 22]. Software [121, 364, 509, 210, 502]. Some [36, 123]. Sons [518]. Sorting [456]. Sound [24]. Sozialforschung [254]. Sozialwissenschaften [100, 60]. specialisés [39]. SPECTRAL [65]. Springer [457, 500]. Springer-Verlag [457]. SPS [402]. SPSS [506, 509, 518, 457, 500, 459, 513, 141, 99, 24, 380, 151, 152, 153, 154, 155, 156, 157, 159, 177, 178, 162, 13, 25, 26, 163, 164, 165, 179, 180, 181, 182, 183, 184, 185, 186, 209, 210, 272, 306, 307, 308, 309, 331, 187, 407, 440, 310, 408, 166, 508, 188, 77, 78, 189, 39, 79, 121, 381, 409, 27, 190, 122, 191, 40, 53, 15, 211, 133, 14, 192, 101, 80, 102, 54, 227, 273, 255, 442, 167, 168, 382, 16, 81, 82, 256, 443, 103, 41, 410, 383, 411, 364, 257]. SPSS [258, 412, 413, 5, 123, 10, 55, 105, 259, 365, 414, 510, 83, 84, 384, 517, 415, 124, 125, 126, 106, 275, 214, 312, 313, 416, 385, 444, 17, 276, 314, 215, 85, 127, 134, 86, 333, 11, 169, 366, 456, 29, 315, 387, 42, 505, 43, 417, 170, 136, 1, 2, 3, 4, 8, 9, 18, 19, 20, 21, 30, 31, 44, 45, 46, 56, 57, 58, 59, 60, 87, 88, 107, 108, 228, 511, 89, 137, 171, 172, 173, 194, 216, 217, 218, 229, 230, 231, 232, 233, 234, 235, 260, 261, 262, 277, 278, 279, 280, 281, 282, 283]. SPSS [284, 285, 286, 316, 334, 335, 336, 337, 338, 339, 340, 341, 342, 343, 367, 368, 369, 370, 371, 372, 373, 374, 388, 389, 390, 391, 392, 393, 418, 419, 420, 421, 422, 423, 448, 449, 32, 33, 61, 62, 63, 64, 65, 66, 67, 109, 110, 111, 112, 113, 114, 115, 116, 117, 22, 68, 69, 34, 47, 70, 90, 118, 12, 6, 519, 520, 72, 521, 522, 479, 236, 375, 71, 119, 139, 513, 394, 424, 219, 220, 226, 48, 141, 49, 50, 93, 94, 95, 96, 128, 97, 74, 98, 129, 203, 142]. SPSS [51, 75, 7, 35, 130, 204, 205, 499, 23, 36, 37, 131, 330, 208, 271, 206, 379, 76, 132, 305, 38, 52, 28, 406, 504, 100, 503, 386, 501, 458]. SPSS-11 [183, 136, 107]. SPSS-6000 [189, 83, 84, 9, 205]. SPSS-GENLIN [511]. SPSS-PC [456, 285]. SPSS-X [179, 180, 184, 185, 186, 272, 259, 171, 216, 279, 176, 197, 200, 237, 270, 291, 300, 301, 303, 304, 317, 319, 465, 495, 496, 497, 203, 210]. SPSS-X^{T M} [163]. SPSS/ONLINE [40, 64, 204, 26, 35, 159]. SPSS/PC [380, 182, 209, 308, 309, 310, 381, 409, 211, 273, 255, 383, 411, 364, 257, 258, 412, 413, 365, 415, 385, 333, 315, 387, 194, 217, 218, 229, 232, 233, 234, 235, 260, 261, 262, 280, 281, 282, 283, 284, 286, 316, 339, 340, 341, 342, 343, 368, 371, 372, 373, 374, 389, 390, 199, 240, 241, 242, 243, 244, 245, 246, 250, 253, 264, 265, 266, 267, 268, 269, 288, 289, 292, 293, 294, 295, 296, 297, 298, 324, 325, 326, 327, 328, 329, 345, 347, 356, 357, 359, 360, 377, 395, 396, 397, 399, 400, 401, 404, 405, 432]. SPSS/PC [433, 450, 462, 463, 466, 468, 474, 475, 482, 485, 486, 492, 498, 424, 219, 410, 278, 367, 429, 386]. SPSS/PCb [369, 370, 394]. SPSS/PCp [416, 388, 247]. SPSS/Pro [198]. SPSS/Pro^{T M} [164]. SPSSON8 [105]. SPSS^{T M} [514]. SPSSX [166, 193, 277, 344, 223]. Stanford [15]. Start [273, 255, 295]. Started [317, 318, 462, 463, 464, 465, 288, 289, 345, 395]. Starting [383, 411]. STAT [208]. State [208]. statements [103]. STATISTICA [505]. Statistical [380, 187, 166, 82, 364, 509, 106, 312, 215, 4, 58, 60, 89, 172, 69, 6, 73, 222, 248, 378, 141, 49, 50, 98, 35, 379, 254, 192, 443, 502, 457, 51, 503, 76, 13, 100, 14, 16, 103, 5, 104, 214, 17, 314, 1, 2, 3, 8, 20, 21, 30, 70, 12, 92, 7, 37]. Statistics [121, 442, 382, 410, 213, 314, 333, 169, 500, 505, 137, 171, 173, 216, 217, 229, 233, 234, 279, 281, 334, 335, 338, 340, 388, 390, 391, 393, 418, 420, 421, 423, 445, 447, 449, 507, 138, 325, 425, 434, 436, 451, 467, 468, 469, 481, 485, 487, 489, 495, 512, 394, 208, 207, 524, 212, 517, 518, 339, 343, 293, 271]. Statistik [60, 100]. Statistik-Programm-System [60, 100]. Statistikprogramm [314]. Statistiksoftware [121]. statistique [39]. statistische [215]. statistischer [254]. Statistisches [314]. Statpak [364]. Steed [458]. Step [442]. Step-by-Step [442]. Stephen [500]. StrateGen [450]. Strednej [28]. strojníckej [28]. Structured [121]. Student [335, 338, 467, 481]. Students [509, 431]. Studentware [284, 285, 368, 369, 370, 371, 372, 373, 374, 381, 409, 283]. Stuttgart [406]. Summary [92, 117].

SunOS [349]. Supplement [10, 249, 271]. support [312, 121]. Supported [314]. Survey [41, 10]. SURVIVAL [66]. Susan [500]. Sybase [306, 307]. Symposium [524]. synopses [177]. Syntax [398, 426, 452, 470]. System [102, 81, 275, 169, 135, 58, 59, 60, 87, 107, 108, 228, 336, 389, 392, 419, 422, 446, 397, 398, 403, 426, 455, 470, 471, 163, 165, 79, 103, 136, 19, 342, 140, 95, 96, 76, 38, 100, 330]. système [448]. systemization [443]. Systems [100, 215, 435, 488, 99, 180, 223, 224, 97]. Systemwide [26].

T [161]. T-tests [161]. Tables

[225, 250, 270, 296, 303, 359, 432, 438, 453, 463, 486, 493, 147, 247, 268, 345, 117]. Taking [212]. tasks [443]. tatbiqi [424]. tatbiqiyat [424]. Teaching [14, 23, 220]. tears [134]. Techniques [442]. terminal [168, 220]. Terminalbruk [27]. terminals [49, 50]. tests [161, 63]. TETRACHORIC [32]. Texas [183]. think [512]. Time [195]. Timesharing [26]. tonggye [416]. Tool [104, 42, 24]. Tour [265]. TRANSLATE [253]. transport [179]. Trends [269, 289, 297, 299, 300, 360, 433, 439]. TriMetrix [399, 400]. trois [39]. Tromsø [97]. ts [206]. tutorial [199, 93, 132]. two [220].

u [100]. U.S. [174, 200, 128]. unbalanced [71]. understand [513]. Understanding [518, 517]. Unicos [272]. UNIVAC [27, 190]. Univariate [504, 510, 517, 518]. Universitetet [97]. University [103, 5]. UNIX [418, 419, 420, 446, 362, 434, 435, 436, 455, 487, 488, 489]. Unterstutzung [121]. Update [77, 125, 4, 8, 9, 31, 44, 45, 46, 59, 88, 239, 322, 350, 494, 35, 205, 13, 163, 189, 84, 313, 57, 267, 329]. úpr [28]. US$32.95 [457]. USA [518]. usage [85, 76]. USDA [106]. Use [256, 228, 394, 99, 5, 123, 56, 493, 49, 50, 128, 74, 132]. User [410, 10, 135, 58, 87, 108, 334, 336, 389, 392, 419, 422, 446, 70, 176, 224, 252, 301, 346, 397, 403, 435, 455, 469, 471, 488, 496, 497, 7, 515, 187, 122, 191, 53, 105, 223, 38]. user-oriented [38]. USERPROC [202, 304]. Users [121, 1, 2, 519, 520, 521, 522, 404, 492, 33, 506]. Using [380, 407, 440, 190, 15, 102, 442, 168, 169, 505, 67, 516, 344, 140, 202, 405, 498, 459, 379, 382, 106, 214, 456, 500, 69, 246, 304, 226, 503, 103, 500]. UT [40, 53]. UT-CDC [53]. utilisation [48]. utilizacion [68, 22]. Utilization [254, 220].

v [28]. V2.0 [281, 286, 242, 293, 298, 359]. V3.0 [308, 327, 329]. V3.1 [329]. value [309]. variables [160, 456]. Variance [167, 366, 144, 146, 161]. VAX [180, 210, 193, 196, 223, 251, 291, 317, 319, 321, 363, 465, 490, 208, 140]. VAX/VMS [223, 317, 319, 321, 363, 490, 291]. väzbu [28]. ved [97]. Verlag [406, 457]. Version [26, 387, 43, 9, 30, 388, 389, 446, 237, 321, 397, 455, 466, 204, 205, 37, 13, 53, 105, 86, 448, 92, 197, 291, 52]. Versionen [375]. Versions [100, 500, 375]. VI [202, 304]. videorecording [143, 144, 145, 146, 147, 148, 149, 158, 159, 160, 161, 103]. vii [457, 500]. VMS [223, 291, 317, 319, 321, 363, 465, 490]. Voelkl [457, 500]. výrociu [28].

W [500]. wa [416]. way [161]. Weber [208]. Webster [394]. Weiss [271]. Weiss/Hassett [271]. who [512]. Wiley [518, 458]. Windows [500, 411, 444, 391, 392, 393, 421, 422, 423, 448, 403, 404, 425, 437, 438, 439, 451, 453, 454, 491, 492, 493, 494, 500, 459]. Wissenschaftler [314]. within [202, 304]. Without [458, 134]. Wittenberg [406].

X [210, 386, 501, 458, 179, 180, 184, 185, 186, 272, 259, 457, 171, 172, 173, 216, 279, 174, 175, 176, 197, 200, 201, 202, 225, 237, 251, 252, 270, 291, 300, 301, 302, 303, 304, 317, 319, 465, 495, 496, 497, 203]. xii [503]. XT [229, 234, 281, 286, 339, 341, 343, 242, 253, 287, 288, 289, 293, 298, 318, 329, 345, 359, 182, 194, 218, 232, 235, 282, 241, 245, 246, 250, 264, 267, 269, 290, 295, 328, 356, 357, 360]. XTAB [29]. XTAB/SPSS [29]. X^{T M} [163]. xvi [500]. xvii [500]. xxi [518].

York [457, 500]. yung [206].

zahlreichen [314]. zost [28].

References

Nie:1968:SSP

[1] Norman H. Nie, Dale H. Bent, and C. Hadlai Hull. SPSS: Statistical Package for the Social Sciences (Provisional Users Manual). SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1968. xi + 425 pp. See [2].

Nie:1969:SSP

[2] Norman H. Nie, Dale J. Bent, and C. Hadlai Hull. SPSS: Statistical Package for the Social Sciences (Provisional Users Manual). SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1969. xi + 425 pp. LCCN ?? First published in [1]; reprinted with the addition of Section XVI in March 1969. Bibliographical footnotes.

Nie:1970:SSP

[3] Norman H. Nie, Dale H. Bent, and C. Hadlai Hull. SPSS: Statistical Package for the Social Sciences. McGraw-Hill, Inc., New York, NY, USA, 1970. xx + 343 pp. LCCN HA33 .N481.

Nie:1971:SSP

[4] Norman H. Nie and C. Hadlai Hull. SPSS: Statistical Package for the Social Sciences: Update Manual. National Opinion Research Center, University of Chicago, Chicago, IL, USA, 1971. various pp.

Green:1972:SSP

[5] Justin Green. SPSS: Statistical Package for the Social Sciences: a guide for use at the University of Iowa, level 3.0. Technical Report 3-30, Laboratory for Political Research, University of Iowa, Iowa City, IA, USA, 1972. 58 pp.

Reynolds:1972:DSS

[6] Henry T. Reynolds and Barbara H. Page. Delaware SPSS: Statistical package for the social sciences. Technical report, University of Delaware, Newark, DE, USA, 1972. 47 pp.

UCD:1972:BBS

[7] University of California, Davis. Institute of Governmental Affairs. Social Sciences Data Services, Davis, CA, USA. Burroughs B6700 SPSS: Statistical Package for the Social Sciences; User’s Manual, 1972. 155 pp.

Nie:1973:SSP

[8] Norman H. Nie. SPSS: Statistical Package for the Social Sciences: Update manual. Technical report, National Opinion Research Center, University of Chicago, Chicago, IL, USA, 1973. various pp.

Nie:1973:SUM

[9] Norman H. Nie. SPSS-6000: Update manual version 5.5. Technical report, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1973. vii + 96 pp.

Hays:1974:UNS

[10] Timothy Paul Hays and Concepcion Shimizu Wilson. A User and Nonuser Survey of Public Libraries in Region G, North Carolina, Explanation of SPSS Program and the Coding of Data: Supplement. Piedmont Triad Council of Governments, Greensboro, NC, USA, 1974. ii + A–H + 101 pp.

Leuthold:1974:IHS

[11] David A. Leuthold, Bob Jakes, and Don McNeil. Introductory handbook to SPSS computer package. Technical report, University of Missouri-Columbia, Dept. of Political Science, Columbia, MO, USA, 1974. 85 pp.

QUBCC:1974:SSP

[12] Queen’s University of Belfast, Computer Centre, Belfast, Northern Ireland. SPSS: Statistical Package for the Social Sciences Introductory Guide, 1974. 23 pp.

Anonymous:1975:SSP

[13] Anonymous. SPSS, Statistical Package for the Social Sciences: [SPSS 6000 version 6.0 update manual]. Technical report, University Computer Center, Minneapolis, MN, USA, 1975.

Charlick:1975:EAS

[14] Robert B. Charlick and Norman H. Nie. Exercises and answers for SPSS mini-manual to accompany self-teaching guide to: Statistical Package for the Social Sciences (SPSS). Technical report, Dept. of Political Science, Cleveland State University, Cleveland, OH, USA, 1975. 73 pp.

CIPSU:1975:USA

[15] Center for Information Processing. Using SPSS at Stanford. Technical report, Stanford Center for Information Processing, Stanford University, Stanford, CA, USA, December 1975. 10 + [16] pp.

Downing:1975:PIS

[16] Bruce B. Downing. A programmed introduction to SPSS: (Statistical Package for the Social Sciences). Technical report, Learning Resources in International Studies, New York, NY, USA, 1975. 35 pp. Revised edition, incorporating version 6 of SPSS.

Klecka:1975:SPS

[17] William R. Klecka, Norman H. Nie, and C. Hadlai Hull. SPSS Primer: Statistical Package for the Social Sciences Primer. McGraw-Hill, Inc., New York, NY, USA, 1975. ISBN 0-07-035023-X. x + 134 pp. LCCN H61 .K551.

Nie:1975:DBS

[18] Norman H. Nie. Differences between SPSS 6.0/CDC6000 and the SPSS manual. Technical report, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1975.

Nie:1975:NDM

[19] Norman H. Nie. NUCC documents and manuals relating to the SPSS system. Technical report, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1975. various pp.

Nie:1975:SSPa

[20] Norman H. Nie. SPSS, Statistical Package for the Social Sciences. McGraw-Hill, Inc., New York, NY, USA, second edition, 1975. ISBN 0-07-046531-2. xxiv + 675 pp. LCCN HA33 .N665 1975.

Nie:1975:SSPb

[21] Norman H. Nie. SPSS: Statistical Package for the Social Sciences. McGraw-Hill, Inc., New York, NY, USA, microfiche edition, 1975. 11 microfiche.

Padua:1975:PEP

[22] Jorge Padua N. Paguete Estadístico para las Ciencias Sociales (SPSS), oferta y condiciones para su utilización e interpretación de resultados. Technical report, Centro de Estudios Sociológicos, Colegio de Mexico, Mexico, DF, Mexico, 1975. 104 pp.

UNTCL:1975:SSC

[23] University of Newcastle upon Tyne. Computing Laboratory. SPSS selfteaching course. Technical report, University of Newcastle upon Tyne, Newcastle upon Tyne, England, 1975. 413 slides.

ALA:1976:SLR

[24] American Library Association. Library Research Round Table. SPSS as a library research tool / library research round table forum IV. [sound recording]. Technical report, American Library Association, Chicago, IL, USA, July 1976. 1 cassette.

Anonymous:1976:HRS

[25] Anonymous. How to run SPSS on the DECsystem-10. Technical report, Computing Facility, University of California, Irvine, Irvine, CA, USA, 1976. 34 pp.

Anonymous:1976:SOV

[26] Anonymous. SPSS/online version 4.0: Systemwide timesharing instructional application, fall 1976. Technical report, Division of Information Systems, California State University and Colleges, Long Beach, CA, USA, 1976.

Brosveet:1976:TMS

[27] Jarle Brosveet. Terminalbruk med SPSS på UNIVAC. NSD rapport nr. 7, Norsk samfunnsvitenskapelig datatjeneste. Norsk samfunnsvitenskapelig datatjeneste, Bergen, Norway, 1976. ISBN 82-7170-011-1. 44 pp. LCCN ??

Hrin:1976:IDR

[28] Milan Hriň and Stredná priemyselná škola strojnícka v. Martine. 25 [i.e. dvadsaťpäť] rokov SPŠS Martin: pamätnica k 25. výročiu strednej priemyselnej školy strojníckej v martine / zost. Milan Hriň a kol; [ob. a väzbu a graf. úpr. navrhol Ján Andel; fot. Ján Grančič]. Technical report, Osveta pre SPŠS??, 1976. 143 pp.

Macbriar:1976:XSD

[29] I. D. Macbriar. An XTAB/SPSS data conversion program. Technical Report CTS WN 26, Centre for Transport Studies, Cranfield Institute of Technology, Cranfield, England, 1976. 10 pp.

Nie:1976:PGV

[30] Norman H. Nie. Programming guide to version 6.0 of Statistical Package for the Social Sciences (SPSS). Data analysis note 2, United States General Accounting Office, Federal Personnel and Compensation Division, Systems Analysis Group, Washington, DC, USA, 1976. 34 pp.

Nie:1976:SRU

[31] Norman H. Nie. SPSS Release 7.0 Update Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1976. 140 pp.

OCSGIT:1976:ST

[32] Office of Computing Services, Georgia Institute of Technology. SPSS TETRA-CHORIC. Technical information memorandum TIM013, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1976.

OCSGIT:1976:SUG

[33] Office of Computing Services, Georgia Institute of Technology. SPSS: users guide to RELIABILITY. Technical information memorandum TIM021, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1976.

Precht:1976:MSR

[34] Dan Precht. MTS/SPSS reference manual. Technical report, Computing Services, University of Alberta, Edmonton, Alberta, Canada, 1976. 102 pp.

UCS:1976:SOS

[35] United Computing Systems, Inc. SPSS/ online: Statistical package for the social sciences: Update manual. Technical report, United Computing Systems, Inc., Kansas City, MO, USA, 1976. [3] + 39 pp.

USGAO:1976:ESS

[36] Anonymous. Explanations of some SPSS procedures. Data analysis note 1, United States General Accounting Office, Federal Personnel and Compensation Division, Systems Analysis Group, Washington, DC, USA, 1976. various pp.

USGAO:1976:PGV

[37] Anonymous. Programming guide to version 6.0 of Statistical Package for the Social Sciences (SPSS). ITS: Data analysis note 2.2, United States General Accounting Office, Systems Analysis Group, Washington, DC, USA, 1976. 34 pp.

Yao:1976:UIS

[38] Ben-Liou Yao. A user-oriented interactive system for SPSS batch program generation. Thesis (m.s.), University of Rhode Island, Kingston, RI, USA, 1976. 55 + [28] pp.

Bernard:1977:CTL

[39] G. Bernard. Comparaison de trois logiciels specialisés pour l’analyse statistique: GENSTAT, BMDP, SPSS. Technical report, Departement de mathématique, Université d’Orsay, Orsay, France, 1977. 15 pp.

CCUTA:1977:SOP

[40] Computation Center, University of Texas at Austin, Austin, TX, USA. SPSS/ONLINE pocket reference list: UT ODC 6000, 1977. 4 pp.

FieldInstitute:1977:SCB

[41] The Field Institute. Survey code book: data file no. F17703 (SPSS). Technical report, The Field Institute, San Francisco, CA, USA, 1977. ca. 200 pp.

Marchant:1977:SLR

[42] Maurice P. Marchant, Nathan M. Smith, and Keith H. Stirling. SPSS as a library research tool. Occasional research paper 1, School of Library and Information Sciences, Brigham Young University, Provo, UT, USA, 1977. 44 pp.

McNett:1977:SMF

[43] Charles W. McNett, Herbert Barry, and George Peter Murdock. SPSS Master File Version of the Ethnographic Atlas. Human Relations Area Files, Inc., New Haven, CT, USA, 1977. 20 pp.

Nie:1977:SBRa

[44] Norman H. Nie and C. Hadlai Hull. SPSS batch release 7.0 update manual. Technical Report 4, University Computer and Center, City University of New York, New York, NY, USA, 1977. v + 83 pp.

Nie:1977:SBRb

[45] Norman H. Nie and C. Hadlai Hull. SPSS Batch Release 7.0 Update Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1977. ISBN ?? 135 pp. LCCN ??

Nie:1977:SBRc

[46] Norman H. Nie and C. Hadlai Hull. SPSS Batch Release 7.0: Update Manual. McGraw-Hill, Inc., New York, NY, USA, March 1977. 141 pp.

Precht:1977:MSR

[47] Dan Precht. MTS/SPSS reference manual. Technical report, Computing Services, University of Alberta, Edmonton, Alberta, Canada, 1977. 116 pp.

Tellier:1977:SUF

[48] Gilles Tellier. SPSS: utilisation des fichiers-SPSS. Technical Report B57-01a, Centre de calcul, Université de Montréal, Montréal, Québec, Canada, 1977. iii + 35 pp.

Thistlethwaite:1977:GUSa

[49] Paul C. Thistlethwaite and Dennis A. Rauch. A guide to the use of SPSS (statistical package for the social sciences) on the CDC interactive terminals. Technical Report 1977-1, Center for Business and Economic Research and Department of Marketing and Finance, Western Illinois University, Macomb, IL, USA, 1977. 48 pp.

Thistlethwaite:1977:GUSb

[50] Paul C. Thistlethwaite and Dennis A. Rauch. A guide to the use of SPSS (statistical package for the social sciences) on the CDC interactive terminals. Technical report, Center for Business and Economic Research, Western Illinois University, Macomb, IL, USA, 1977. iv + 79 pp.

UCCCUNY:1977:MSA

[51] University Computer and City University of New York Center. Mathematical and statistical applications: SPSS. Technical Report 4, University Computer and Center, City University of New York, New York, NY, USA, 1977. i + 33 pp.

Zaharris-Esgro:1977:BSG

[52] Rhonda Zaharris-Esgro, Daniel W. Edwards, and Kathy Miyata. Burroughs 6700 SPSS guide: version 6. Technical report, University of California, Davis, Davis, CA, USA, 1977. 77 pp.

CCUTA:1978:SUR

[53] Computation Center, University of Texas at Austin. SPSS user’s reference list: version 6.5 UT-CDC 6000. Technical Report CCRL-3, Computation Center, University of Texas at Austin, Austin, TX, USA, 1978. 14 pp.

Clements:1978:SPG

[54] Ellen Clements. SPSS Pocket Guide: Release 7. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1978. 45 pp.

Hays:1978:SO

[55] Mark Hays and Jonan Mills. SPSS online 7.0. Technical report, California State University and Colleges, Long Beach, CA, USA, 1978. 42 pp.

Nie:1978:NUS

[56] Norman H. Nie. Notes on the use of SPSS, release 7, for Honeywell 6000 and series 60, level 66, computers, April 1978. Technical report, Institute for Social and Environmental Studies, University of Kansas, Lawrence, KS, USA, 1978. 22 pp.

Nie:1978:SBR

[57] Norman H. Nie and C. SPSS batch release 7.0 update manual, march 1977. Technical report, University of Kansas, Lawrence, KS, USA, 1978. ca. 200 pp.

Nie:1978:SSC

[58] Norman H. Nie and C. Hadlai Hull. SCSS: SPSS Conversational Statistical System: Release 2.0, Preliminary User’s Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1978. v + 168 pp.

Nie:1978:SSI

[59] Norman H. Nie and C. Hadlai Hull. SPSS: Batch System Release 8.0 Update Manual, December, 1978. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1978. iii + 112 pp.

Nie:1978:SSS

[60] Norman H. Nie. SPSS 7, Statistik-Programm-System für die Sozialwissenschaften [English: SPSS 7, Statistical Package for the Social Sciences]. G. Fischer, Stuttgart, Germany, second edition, 1978. 276 pp.

OCSGIT:1978:SJ

[61] Office of Computing Services, Georgia Institute of Technology. SPSS JFAC-TOR. Technical information memorandum TIM020, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:SN

[62] Office of Computing Services, Georgia Institute of Technology. SPSS NONLIN-EAR. Technical information memorandum TIM011, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:SNT

[63] Office of Computing Services, Georgia Institute of Technology. SPSS NPAR TESTS. Technical information memorandum TIM012, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:SO

[64] Office of Computing Services, Georgia Institute of Technology. SPSS/ ONLINE. Technical information memorandum TIM005, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:SSa

[65] Office of Computing Services, Georgia Institute of Technology. SPSS SPEC-TRAL. Technical information memorandum TIM024, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:SSb

[66] Office of Computing Services, Georgia Institute of Technology. SPSS SUR-VIVAL. Technical information memorandum TIM023, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

OCSGIT:1978:UFS

[67] Office of Computing Services, Georgia Institute of Technology. Using FOUBAR with SPSS. Technical information memorandum TIM026, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1978.

Padua:1978:PEP

[68] Jorge Padua N. Paquete estadistico para las ciencias sociales (SPSS): oferta y condiciones para su utilizacion e interpretacion de resultados. Technical Report 12, Centro de Estudios Sociologicos, Colegio de Mexico, Mexico, DF, Mexico, 1978. v + 104 pp.

Pooler:1978:DMA

[69] William S. Pooler and Susan Thomas. Data management and analysis using SPSS [statistical package for the social sciences]. Technical Report PS-16, Policy Studies Associates, Crotonon-Hudson, NY, USA, 1978. vi + 65 + [2] pp.

Prime:1978:PPD

[70] Prime Computer, Inc., Framingham, MA, USA. Prime Preliminary Documentation Release: PDR3173 User’s Guide, Statistical Package for the Social Sciences (SPSS), 1978. 157 pp.

Searle:1978:ACO

[71] S. R. (Shayle R.) Searle and Ronald Paul Kershner. Annotated computer output for analyses of unbalanced data: SPSS ANOVA. Biometrics Unit Mimeo series BU-660-M, Biometrics Unit, Cornell University, Ithaca, NY, USA, 1978. 34 pp.

SPSS:1978:SIN

[72] SPSS Inc. Newsletter, 1978. ISSN 0160-7669, 0160-7669. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA.

SPSS:1978:SSA

[73] SPSS, Inc. SPSS, Statistical Algorithms: Release 8A, Preliminary Publication. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1978. x + 162 pp.

Trotter:1978:DEU

[74] William Trotter, Jr. Debugging and efficient use of SPSS. Computer note CN-22, Computer Science Center, University of Maryland, College Park, MD, USA, 1978. 66 pp.

UCCUA:1978:SQR

[75] University Computer Center, University of Arizona. SPSS quick reference. Technical Report UIB-11, University Computing Center, University of Arizona, Tucson, AZ, USA, March 1978. 31 pp.

Wolf:1978:ISC

[76] Thomas E. Wolf. An interactive, selfdocumenting computer interface system to aid usage of the SPSS statistical package [microform]. PhD thesis, University of Pittsburgh, Pittsburgh, PA, USA, 1978. v + 187 pp.

Beadle:1979:SUN

[77] ViAnn Beadle, C. Hadlai Hull, and Norman H. Nie. SPSS Update: New Procedures and Facilities for Releases 7 and 8. McGraw-Hill, Inc., New York, NY, USA, 1979. ISBN 0-07-046534-7. xv + 238 pp. LCCN HA33 .S17.

Benson:1979:SRF

[78] Betty Benson and Mina Hohlen. SPSS report: flexible report generator. Technical Report 456, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1979. 56 pp.

Bloomberg:1979:ISS

[79] David J. Bloomberg, Paul C. Thistlethwaite, and Belle M. Zimmerly. Interactive SPSS: a system file approach, application to advertising. Technical Report 1979-5, Center for Business and Economic Research, Western Illinois University, Macomb, IL, USA, 1979. 1 + 22 pp.

Cleary:1979:SHS

[80] Michael J. Cleary and Robert T. Amsden. The SPSS Helper. Advocate Publishing Group, Reynddsburg, OH, USA, preliminary edition, 1979. ISBN 0-89894-015-X. 88 pp. LCCN ??

Drone:1979:ISS

[81] Sam G. Drone and Paul C. Thistlethwaite. Interactive SPSS: A system file approach, application to finance. Technical Report 1979-4, Center for Business and Economic Research, Western Illinois University, Macomb, IL, USA, 1979. [6] + 22 + [2] pp.

Ebeling:1979:ASS

[82] Jon Ebeling and Michael King. Application of SPSS [statistical package for the social sciences] to data archives available in the CSUC. Technical Report 79-1, School of Behavioral and Social Sciences, California State University, Chico, CA, USA, 1979. 30 pp.

Hohlen:1979:SRI

[83] Mina Hohlen. SPSS-6000 release 8.0 implementation bulletin. Technical report, Northwestern University, Vogelback Computing Center, Evanston, IL, USA, 1979. 1 on–line manual + 1 cover sheet (2 p.) and 1 printout copy of manual (13 p.).

Hohlen:1979:SRU

[84] Mina Hohlen, C. Hadlai Hull, and Norman H. Nie. SPSS-6000 release 8.0 update manual. Technical report, Northwestern University, Vogelback Computing Center, Evanston, IL, USA, 1979. 1 on–line manual + 1 cover sheet and 1 printout copy of manual.

Laforge:1979:AMS

[85] Hubert Laforge. Analyse multivariée: en sciences sociales et biologiques: présentation matricielle et usage des programmes d’ordinateur BMD, BMDP et SPSS. Presses de l’Université Laval, Québec, Québec, Canada, 1979. vi + 450 pp. LCCN ??

Lee:1979:SJS

[86] Chung Lee. SPSS JCL: SPSS version 7.0, job control language. Technical report, Instructional Support Group, Division of Information Systems California State University and Colleges, Long Beach, CA, USA, 1979. 69 pp.

Nie:1979:SSI

[87] Norman H. Nie and C. Hadlai Hull. SCSS: The SPSS Inc. Conversational System: Release 3.1, Preliminary User’s Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1979. v + 188 pp.

Nie:1979:SUN

[88] Norman H. Nie and ViAnn. Beadle. SPSS Update: New Procedure and Facilities for Releases 7 and 8. McGraw-Hill, Inc., New York, NY, USA, 1979. ISBN 0-07-046534-7 (paperback). x + 172 pp. LCCN HA33 .S17.

Norusis:1979:SSA

[89] Marija J. Norušis. SPSS, Statistical Algorithms: Release 8.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1979. xi + 166 pp. LCCN HA33 .S15, HA32 .N67 1979. See [172].

Prime:1979:PSP

[90] Prime Computer, Inc. Prime: The SPSS Programmer’s Guide: PDR3173. Prime Computer, Inc., Framingham. MA, USA, 1979. 155 pp. Revision A.

SPSS:1979:SPG

[91] SPSS, Inc. SPSS Pocket Guide: Release 8. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1979. 50 pp.

SPSS:1979:SSP

[92] SPSS, Inc. SPSS: Statistical Package for the Social Sciences, for the Digital Equipment Corporation PDP-11: Extended Command Summary, September 1978, version 2.2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1979. 47 pp.

Thistlethwaite:1979:BIS

[93] Paul C. Thistlethwaite. Batch and interactive SPSS: a self-paced tutorial. Technical report, Western Illinois University, Center for Business and Economic Research, Macomb, IL, USA, 1979. various pp.

Thistlethwaite:1979:ISA

[94] Paul C. Thistlethwaite. Interactive SPSS: application to recreation. Technical Report 1979-18, Center for Business and Economic Research, College of Business, Western Illinois University, Macomb, IL, USA, 1979. 19 pp.

Thistlethwaite:1979:ISSa

[95] Paul C. Thistlethwaite. Interactive SPSS: a system file approach, application to retailing. Technical Report 1979-3, Center for Business and Economic Research, Western Illinois University, Macomb, IL, USA, 1979. ii + 16 pp.

Thistlethwaite:1979:ISSb

[96] Paul C. Thistlethwaite. Interactive SPSS: a system file approach. Technical report, Center for Business and Economic Research, Western Illinois University, Macomb, IL, USA, February 1979. ii + 45 + [6] pp.

Trollvik:1979:HSV

[97] Jan Trollvik. Håndbok i SPSS ved Cyber computer systems, Universitetet i Tromsø. Number 25 in NSD rapport. Norsk samfunnsvitenskapelig datatjeneste, Bergen, Norway, 1979. ISBN 82-7170-033-2. 92 pp. LCCN ??

UACC:1979:JCL

[98] University of Akron Computer Center. Job control language for SPSS (statistical package for the social sciences). Technical report, University of Akron Computer Center, Akron, OH, USA, July 1979. 24 + [12] pp.

ACCUK:1980:NUS

[99] Academic Computer Center, University of Kansas. Notes on the use of SPSS release 8.1 for Honeywell level 66 and series 6000 GCOS systems. Technical report, University of Kansas, Lawrence, KS, USA, 1980. ii + 34 pp.

Beutel:1980:SSN

[100] Peter Beutel. Statistik-Programm-System fur die Sozialwissenschaften Nach Norman H. Nie, SPSS 8: e. Beschreibung d. Programmversionen 6, 7 u. 8 [English: Statistical Programming Systems for the Social Sciences for Norman H. Nie, SPSS 8: A Description of Program Versions 6, 7, and 8]. G. Fischer, Stuttgart, Germany, third edition, 1980. 300 pp.

Christiansen:1980:BSS

[101] John B. Christiansen. Basic skills in SPSS manual. Gallaudet College, Washington, DC, USA, 1980. xii + 384 + [55] pp.

Cleary:1980:DAH

[102] Michael J. Cleary and Robert T. Amsden. A Data Analysis Handbook Using the SPSS System. Advocate Publishing Group, Reynddsburg, OH, USA, 1980. ISBN 0-89894-015-X. viii + 149 pp. LCCN HA33 .C62.

Fetcher:1980:SVS

[103] Lincoln Fetcher, Elizabeth Hinkley, University of Minnesota. Computer Center, and University of Minnesota. Media Resources. SPSS [videorecording] / a University Computer Center production; directed by Lincoln Fetcher. SPSS, Statistical Package for the Social Sciences. Basic commands for a simple job. Data manipulation statements. Using SPSS system files. Technical report, University Computer Center, University Media Resources, Minneapolis, MN, USA, 1980. 3 videocassettes (145 min.).

Gutwein:1980:SPS

[104] John Gutwein. Statistical Package for the Social Sciences: a tool for planners. Thesis (m.u.p.), Dept. of Urban Planning, California State Polytechnic University, Pomona, Pomona, CA, USA, 1980. xii + 152 + [91] pp.

Hays:1980:SSO

[105] Mark Hays, Jonah Mills, and Gary A.

Jones. SPSSON8: SPSS online version 8 user’s guide. Technical report, Instructional Support Group, CSUC, Los Angeles, CA, USA, 1980. 1 on–line manual + 1 cover sheet (2 p.) and 1 printout copy.

Ingwersen:1980:SAU

[106] James B. Ingwersen. Statistical analysis using SPSS at the USDA–Fort Collins Computer Center. WSDG application document WSDG-AD-00002, Watershed Systems Development Group, USDA Forest Service, Fort Collins, CO, USA, December 1980. 8 + [10] pp.

Nie:1980:SSB

[107] Norman H. Nie, C. Hadlai Hull, and Graham Kimble. SPSS-11, the SPSS Batch System for DEC PDP-11. McGraw-Hill, Inc., New York, NY, USA, 1980. ISBN 0-07-046537-1. viii + 165 pp. LCCN HA 29 S14 1980. See [136].

Nie:1980:SUG

[108] Norman H. Nie. SPSS, a User’s Guide to the SPSS Conversational System. McGraw-Hill, Inc., New York, NY, USA, 1980. ISBN 0-07-046533-9. xxv + 595 pp. LCCN H61.3 .S17.

OCSGIT:1980:SA

[109] Office of Computing Services, Georgia Institute of Technology. SPSS ANOVA. Technical information memorandum TIM015.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SG

[110] Office of Computing Services, Georgia Institute of Technology. SPSS G3SLS. Technical information memorandum TIM017.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SM

[111] Office of Computing Services, Georgia Institute of Technology. SPSS MANOVA. Technical information memorandum TIM007.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SMR

[112] Office of Computing Services, Georgia Institute of Technology. SPSS MULT RESPONSE. Technical information memorandum TIM022.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SP

[113] Office of Computing Services, Georgia Institute of Technology. SPSS PLOT. Technical information memorandum TIM018.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SRa

[114] Office of Computing Services, Georgia Institute of Technology. SPSS REGRESSION. Technical information memorandum TIM009.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SRb

[115] Office of Computing Services, Georgia Institute of Technology. SPSS RELIA-BILITY. Technical information memorandum TIM010.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SRc

[116] Office of Computing Services, Georgia Institute of Technology. SPSS RE-PORT. Technical information memorandum TIM027, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

OCSGIT:1980:SST

[117] Office of Computing Services, Georgia Institute of Technology. SPSS SUM-MARY TABLES. Technical information memorandum TIM016.A, Office of Computing Services, Georgia Institute of Technology, Atlanta, GA, USA, 1980.

Prime:1980:SPC

[118] Prime Computer, Inc., Framingham, MA, USA. SPSS on Prime Computers, 1980. v + 24 pp.

Searle:1980:SA

[119] Shayle R. Searle and R. P. Kershner. SPSS ANOVA. Biometrics Unit Mimeo series BU-660-M, Biometrics Unit, Cornell University, Ithaca, NY, USA, 1980. 34 pp.

SPSS:1980:K

[120] Keywords, 1980. ISSN 0197-7342. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA.

Bodendorf:1981:UAS

[121] Freimut Bodendorf. Unterstutzung der Anwender von Statistiksoftware: ein Methodenbankrahmen um das Programmpaket SPSS [English: Support for users of statistics software: A structured outline about the program package SPSS]. Thesis, Friedrich-Alexander-Universität, Erlangen-Nürberg, Germany, 1981. vii + 157 pp.

CCIC:1981:SUM

[122] Computer Center, Iona College, New Rochelle, NY, USA. SPSS under MU-SIC: a user’s guide, 1981. 101 pp. Version 2.0.

Greenberg:1981:CSA

[123] Edward A. Greenberg, William Max Lewis, and Bruce R. Lewis. Comparison of some available packages for use in research data management: [SIR, SAS, and SPSS]. Technical report, ACM, 1981. 7 pp.

Hull:1981:SPG

[124] C. Hadlai Hull and Norman H.

Nie. SPSS Pocket Guide: Release 9. McGraw-Hill, Inc., New York, NY, USA, 1981. ISBN 0-07-046543-6. 45 pp. LCCN HA32 .H84.

Hull:1981:SUN

[125] C. Hadlai Hull and Norman H. Nie.

SPSS Update 7–9: New Procedures and Facilities for Releases 7-9. McGraw-Hill, Inc., New York, NY, USA, 1981. ISBN 0-07-046542-8 (paperback). xxi + 402 pp. LCCN HA33 .S16 1981.

Humphrey:1981:MSS

[126] C. Humphrey. MTS/SPSS. Technical Report R12.1281, Computing Services, University of Alberta, Edmonton, Alberta, Canada, 1981. 158 pp.

Laforge:1981:AMP

[127] Hubert Laforge. Analyse multivariée:

pour les sciences sociales et biologiques avec applications des logiciels BMD, BMDP, SPSS, SAS. Etudes vivantes, Saint-Lambert, Québec, 1981. ISBN 2-7607-0075-5. xiii + 317 pp. LCCN ??

Tompkins:1981:USE

[128] Linda Tompkins and Clevie E. Gladney. The use of SPSS in the Economic Development Division of the U.S. Department of Agriculture. ESS staff report AGESS810528, Data Services Center, Economics and Statistics Service, U.S. Dept. of Agriculture, Washington, DC, USA, 1981. various pp.

UCB:1981:SUC

[129] Anonymous. SPSS under CMS. Technical report, Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1981. 7 pp.

UGCS:1981:SSU

[130] Anonymous. SPSS on the 2976. Technical report, University of Glasgow, Computing Service, Glasgow G12 8QQ, Scotland, 1981. i + 23 pp.

VCCNWU:1981:SCP

[131] Vogelback Computing Center, Northwestern University. SPSS computer programs series. Technical report, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1981.

Wolf:1981:DCA

[132] Melanie Joy Wolf. Development of a computer assisted tutorial in the use of SPSS for application to health service agencies. Master’s thesis, University of Texas Health Science Center at Houston, School of Public Health, Houston, TX, USA, 1981. 24 pp.

CSUCB:1982:SCG

[133] Computing Services, University of California, Berkeley. SPSS conversion guide: going from the CDC 6400 to the IBM 4341. Technical report, Computing Services, University of California, Berkeley, Berkeley, CA, USA, August 1982. 12 pp.

Lamp:1982:DSM

[134] E. Joseph Lamp and Stephen F. Steele.

Doing SPSS without many tears. Ginn Custom Publishing, Lexington, MA, USA, 1982. 54 pp.

Ling:1982:IUG

[135] Robert F. Ling and Harry V. Roberts.

IDA, a User’s Guide to the IDA Interactive Data Analysis and Forecasting System. The SPSS series in data analysis. Scientific Press, Palo Alto, CA, USA, 1982. ISBN 0-07-037906-8 (paperback). various pp. LCCN QA76.8.I2633 .L5 1982.

Morrison:1982:SSB

[136] Nancy K. Morrison. SPSS-11, the SPSS batch system for the DEC PDP-11. McGraw-Hill, Inc., New York, NY, USA, second edition, 1982. ISBN 0-07-046546-0 (paperback), 0-918469-28-7 (paperback). xvi + 312 pp. LCCN HA32 .M67 1982.

Norusis:1982:SIG

[137] Marija J. Norušis, Norman H. Nie, and C. Hadlai Hull. SPSS Introductory Guide: Basic Statistics and Operations. McGraw-Hill, Inc., New York, NY, USA, 1982. ISBN 0-07-047528-8. xiv + 173 pp. LCCN HA32 .N66.

Roberts:1982:CSI

[138] Harry V. Roberts and Robert F. Ling.

Conversational Statistics with IDA: An Introduction to Data Analysis and Regression. The SPSS series in data analysis. Scientific Press, Palo Alto, CA, USA, 1982. ISBN 0-07-053137-4. 648 pp. LCCN HD30.23.R6 1982.

Searle:1982:ACO

[139] S. R. (Shayle R.) Searle, Walter Theodore Federer, and G. F. S. Hudson. Annotated computer output for covariance (ACO: COV — SPSS). Biometrics Unit Mimeo series BU-784-M, Biometrics Unit, Cornell University, Ithaca, NY, USA, July 1982.

SPSS:1982:USB

[140] SPSS, Inc. Using SPSS batch system, release 9.1 on the Digital Vax. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1982. 28 pp.

Thaller:1982:NDH

[141] Manfred Thaller. Numerische Datenverarbeitung für Historiker: eine praxisorientierte Einführung in die Quantitative Arbeitsmethode und in SPSS (Statistical Package for the Social Sciences) [English: Numerical Data Processing for Historians: a Practical Introduction in Quantitative Methods and in SPSS]. Materialien zur historischen Sozialwissenschaft; 1. Hermann Böhlaus, Wien, Austria, 1982. ISBN 3-205-07074-7. xii + 166 pp. LCCN HA 29.5 G4 T45 1982.

UCBCS:1982:SCG

[142] Computing Services, University of California, Berkeley. SPSS conversion guide: going from the CDC 6400 to the IBM 4341. Technical Report CDC 14.2.6, Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1982. 12 pp.

Ahlgren:1983:CDV

[143] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Cleaning up data [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (42 min.).

Ahlgren:1983:FAV

[144] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Factorial analysis of variance and covariance [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (39 min.).

Ahlgren:1983:IVS

[145] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Introduction [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (18 min.).

Ahlgren:1983:MAV

[146] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Multivariate analysis of variance and covariance [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (73 min.).

Ahlgren:1983:MCT

[147] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Multidimensional contingency tables [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (72 min.).

Ahlgren:1983:MFA

[148] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. More factor analysis [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (67 min.).

Ahlgren:1983:MMR

[149] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. More multiple regression [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (41 min.).

Ahlgren:1983:MR

[150] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Multiple regression. Practical data analysis with SPSS; no. 8, pt. A-B. University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (72 min.).

Ahlgren:1983:PDAa

[151] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (67 min.).

Ahlgren:1983:PDAb

[152] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (42 min.) plus guide.

Ahlgren:1983:PDAc

[153] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (18 min.) plus guide.

Ahlgren:1983:PDAd

[154] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (58 min.) plus guide.

Ahlgren:1983:PDAe

[155] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (38 min.) plus guide.

Ahlgren:1983:PDAf

[156] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (39 min.).

Ahlgren:1983:PDAg

[157] Andrew Ahlgren. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (73 min.).

Ahlgren:1983:RRV

[158] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Reporting results [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 2 videocassettes (74 min.).

Ahlgren:1983:SOV

[159] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. SPSS/online [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (35 min.).

Ahlgren:1983:SVV

[160] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. Shaping up variables [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (59 min.).

Ahlgren:1983:TOA

[161] Andrew Ahlgren, Patricia Heller, and Nancy Mosier. T-tests and one-way analysis of variance [videorecording]. Technical report, Media Resources, University of Minnesota, Minneapolis, MN, USA, 1983. 1 videocassette (38 min.).

Allen:1983:HSO

[162] William R. Allen. A handbook for SPSS online auto-mode. Technical Report 83-5, School of Behavioral and Social Sciences, California State University, Chico, CA, USA, 1983. 19 pp.

Anonymous:1983:SIA

[163] Anonymous. SPSS Inc. announces new LISREL procedure in first update of SPSS-X^{T M} information analysis system. Computational Statistics & Data Analysis, 1(??):278–279, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901020.

Anonymous:1983:SII

[164] Anonymous. SPSS Inc. introduces SPSS/Pro^{T M}. Computational Statistics & Data Analysis, 1(??):279, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947383901032.

Anonymous:1983:STS

[165] Anonymous. The SPSS^{X} ^{T M} system. Computational Statistics & Data Analysis, 1(??):63, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900695.

Barcikowski:1983:CPR

[166] Robert S. Barcikowski. Computer Packages and Research Design: With Annotations of Input and Output from the BMDP, SAS, SPSS, and SPSSX Statistical Packages. University Press of America, Lanham, MD, USA, 1983. ISBN 0-8191-3494-5 (v. 1: paperback), 0-8191-3495-3 (v. 2: paperback), 0-8191-3496-1 (v. 3: paperback). xiv + 781 pp. LCCN QA276.4.C576 1983.

DeBoer:1983:AVC

[167] Donald P. DeBoer, Frederick T. Evers, and Douglas Garfield Link. Analysis of Variance: Comparative Examples, BAL-ANOVA, BMDP and SPSS. Social Science Computing Laboratory, University of Western Ontario, London, Ontario, Canada, 1983. ISBN 0-7714-0427-1. various pp. LCCN ??

Doerr:1983:UST

[168] Helen M. Doerr. Using SPSS through a terminal. Technical report, State University of New York College at Cortland, Cortland, NY, USA, 1983. 14 pp.

Levine:1983:USB

[169] David M. Levine, Mark L. Berenson, and David Stephan. Using the SPSS Batch System with Basic Business Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1983. ISBN 0-13-058719-2 (paperback). vi + 69 pp. LCCN ??

Millikan:1983:PPG

[170] Jane A. Millikan. Perkin-Elmer primer:

a guide to EDIT and SPSS. McLean Hospital Evaluative Service Unit Reports 54, McLean Hospital Evaluative Service Unit, Belmont, MA, USA, 1983. 17 pp.

Norusis:1983:SIS

[171] Marija J. Norušis. SPSS-X Introductory Statistics Guide for SPSS-X Release 3. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1983.

Norusis:1983:SSA

[172] Marija J. Norušis. SPSS X, Statistical Algorithms. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1983. xiii + 194 pp. LCCN HA32 .S65 1983.

Norusis:1983:SXI

[173] Marija J. Norušis. SPSS X Introductory Statistics Guide. McGraw-Hill, Inc. and SPSS, Inc., New York, NY, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1983. ISBN 0-07-046549-5 (paperback), 0-918469-20-1 (paperback). xii + 276 pp. LCCN HA32 .N681 1983.

SPSS:1983:SPU

[174] SPSS, Inc. SPSS X Processing of U.S. Census Data. SPSS Inc. reports, No. 4. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1983. ISBN 0-918469-02-3 (paperback). 79 pp. LCCN CM.1.1.03.

SPSS:1983:SRN

[175] SPSS, Inc. SPSS X Release 2, New Procedures and Facilities. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1983. iii + 104 pp.

SPSS:1983:SUG

[176] SPSS, Inc. SPSS-X User’s Guide.

McGraw-Hill, Inc., New York, NY, USA, 1983. ISBN 0-07-046550-9 (paperback). xxv + 806 pp. LCCN HA32 .S67 1983.

Ahlgren:1984:PDAh

[177] Andrew Ahlgren, Patricia Heller, and Nancy R. Mosier. Practical data analysis with SPSS: synopses and annotations. University of Minnesota, Minneapolis, MN, USA, 1984. 137 pp.

Ahlgren:1984:PDAi

[178] Andrew Ahlgren, Patricia Heller, and Nancy R. Mosier. Practical data analysis with SPSS. University of Minnesota, Minneapolis, MN, USA, 1984. Videocassettes (625 min.).

Anonymous:1984:IEF

[179] Anonymous. IMPORT/EXPORT facility in SPSS-X to transport data and dictionary files. Computational Statistics & Data Analysis, 2 (1):104, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900446.

Anonymous:1984:NRS

[180] Anonymous. New release of SPSS-X now available on DEC VAX systems. Computational Statistics & Data Analysis, 2(1):108–110, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900495.

Anonymous:1984:SIA

[181] Anonymous. SPSS Inc. announces new consulting group. Computational Statistics & Data Analysis, 2 (1):100, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900410.

Anonymous:1984:SPI

[182] Anonymous. SPSS/PC for the IBM PC XT. Computational Statistics & Data Analysis, 2(1):103, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900434.

Anonymous:1984:STI

[183] Anonymous. SPSS-11 for the Texas instruments 990. Computational Statistics & Data Analysis, 2(2): 179, August 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900057.

Anonymous:1984:SXAa

[184] Anonymous. SPSS-X^{T M} available on Honeywell GCOS. Computational Statistics & Data Analysis, 2 (1):110, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900501.

Anonymous:1984:SXAb

[185] Anonymous. SPSS-X available on prime computers. Computational Statistics & Data Analysis, 2(2):179– 180, August 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900069.

Anonymous:1984:SXH

[186] Anonymous. SPSS-X on Hewlett– Packard. Computational Statistics & Data Analysis, 2(1):108, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/0167947384900483.

Arigo:1984:CCP

[187] R. Arigo, S. E. Howe, and T. Webb.

Climatic calibration of pollen data: a user’s guide for the applicable computer programs in the statistical package for social scientists (SPSS). Technical report, Division of Radiation Programs and Earth Sciences, Office of Nuclear Regulatory Research, U.S. Nuclear Regulatory Commission, Washington, DC, USA, June 1984. vii + 36 pp.

Bauer:1984:DSS

[188] Dipl.-Sozialwirt Felix Bauer. Datenanalyse mit SPSS. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-13269-4 (New York). ix + 275 pp. LCCN HA32 .B38 1984.

Benson:1984:SUM

[189] Betty Benson and Mina Hohlen. SPSS-6000 update manual. Technical report, Vogelback Computing Center, Northwestern University, Evanston, IL, USA, 1984. iv + 86 pp. Revised for version 9.0.

Butler:1984:USU

[190] Bertha Butler and Mitchell Karpman. Using SPSS on the UNIVAC 1100. Computer note CN-22.2, Computer Science Center Publications, University of Maryland, College Park, MD, USA, September 1984. 20 pp.

CCIC:1984:SUM

[191] Computer Center, Iona College, New Rochelle, NY, USA. SPSS under MU-SIC: a user’s guide, September 1984. 85 pp. Version 3.0.

Chongzi:1984:ESA

[192] Lei Chongzi. The exploration of SPSS and the applications of multivariate statistical analysis. Technical report, Management School of SJTU, ??, 1984. 7 pp.

LarsonAlbers:1984:ADS

[193] Cathy Larson-Albers. Accessing DATA-TRIEVE in SPSSX on the digital VAX. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. 15 pp.

Norusis:1984:SPI

[194] Marija J. Norušis. SPSS/PC for the IBM PC/XT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. ISBN 0-918469-00-7 (paperback). 624 pp. LCCN HA32 .N665 1984.

Roberts:1984:TSF

[195] Harry V. Roberts. Time Series and Forecasting with IDA. The SPSS series in data analysis. Scientific Press, Palo Alto, CA, USA, 1984. ISBN 0-07-053136-6 (paperback). 500 pp. LCCN HD30.27.R6 1984.

SPSS:1984:ADS

[196] SPSS, Inc. Accessing DATATRIEVE in SPSS on the DIGITAL VAX. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984.

SPSS:1984:SDF

[197] SPSS, Inc. SPSS-X data file formats:

IBM 370 version. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. iv + 28 + [1] pp.

SPSS:1984:SDP

[198] SPSS, Inc. SPSS/Pro for DEC Professional 350. SPSS, Inc. and McGraw-Hill, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and New York, NY, USA, 1984. ISBN 0-07-046556-8 (paperback). xii + 282 pp. LCCN HA32 .S63 1984.

SPSS:1984:SPT

[199] SPSS, Inc. SPSS/PC tutorial. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984.

SPSS:1984:SPU

[200] SPSS, Inc. SPSS-X processing of U.S.

Census data. SPSS Inc. reports no. 4. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. 79 pp.

SPSS:1984:SXB

[201] SPSS, Inc. SPSS X Basics. McGraw-Hill, Inc. and SPSS, Inc., New York, NY, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. ISBN 0-07-060524-6 (paperback) and 0-918469-22-8. ix + 214 pp. LCCN HA32 .S64

SPSS:1984:ULU

[202] SPSS, Inc. USERPROC LISREL: Using LISREL VI within SPSS X. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984. 34 pp.

UCB:1984:SUC

[203] University of California, Berkeley, Computing Services, Berkeley, CA, USA. SPSS-X under CMS, 1984. 8 pp.

UMCC:1984:SOV

[204] University of Minnesota, Computer Center. SPSS/ONLINE version 9.0 SPSS ONLINE version 9.0. Technical report, University of Minnesota, University Computer Center, Minneapolis, MN, USA, 1984. iv + 59 pp.

UMCC:1984:SVU

[205] University of Minnesota, Computer Center. SPSS-6000 version 9.0: Update manual. Technical report, University of Minnesota, University Computer Center, Minneapolis, MN, USA, 1984. v + 81 pp.

Weng:1984:SSY

[206] Shu yüan Weng. SPSS shih yung shou tse. Wu nan tu shu chu pan kung ssu, Min kuo 73, Tai-pei shih, 880-03 chu pan edition, 1984. 502 pp.

Wright:1984:CSE

[207] Benjamin Drake Wright and Patrick L. Mayers. Conversational Statistics in Education and Psychology with IDA. The SPSS series in data analysis. Scientific Press, Palo Alto, CA, USA, 1984. ISBN 0-07-072081-9 (paperback). 237 pp. LCCN LB 2846 W75 1984.

WSUCS:1984:SWS

[208] Weber State University Computing Services. Statistics on the Weber State VAX: SPSS and STAT. Technical Report 2, Weber State University, Ogden, UT, USA, 1984. 60 pp.

Anonymous:1985:MMC

[209] Anonymous. Micro to mainframe communications with SPSS/PC and Kermit. Computational Statistics & Data Analysis, 2(4):337, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900076.

Anonymous:1985:SXS

[210] Anonymous. SPSS-x software for VAX.

Computational Statistics & Data Analysis, 3(??):72–73, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900623.

CSPUCC:1985:CPC

[211] California State Polytechnic University, Computer Center, Pomona, CA, USA. Cal Poly Computer Lab SPSS/PC, 1985. 10 program files (IBM PC) on 1 computer disk.

Fridlund:1985:TNB

[212] Alan J. Fridlund. Taking the numbers bull by the horns: four statistics packages provide a range of power and features: [review]. InfoWorld, ??(??):42– 50, February 11 1985. CODEN IN-WODU. ISSN 0199-6649.

Jendrek:1985:TMS

[213] Margaret Platt Jendrek. Through the Maze: Statistics with Computer Applications. Wadsworth Publishing Company, Belmont, CA, USA, 1985. ISBN 0-534-03921-9. x + 218 pp. LCCN HA32 .J46 1985.

Kellough:1985:SPS

[214] Patrick Henry Kellough. Statistical Package for the Social Sciences Bibliography: Selected SPSS Manuals and Research Articles using SPSS. Public administration series–bibliography P 1671. Vance Bibliographies, Monticello, IL, USA, April 1985. ISBN 0-89028-381-8 (paperback). ISSN 0193-970X. 8 pp. LCCN Z 7551 K38 1985.

Kuffner:1985:DSA

[215] Helmuth Kuffner and Reinhard Wittenberg. Datenanalysessyteme fur statistische Auswertungen: e. Einführung in SPSS, BMDP und SAS [English: Data Analysis Systems for Statistical Evaluations: An Introduction to SPSS, BMDP, and SAS]. G. Fischer, Stuttgart, Germany, 1985. 292 pp.

Norusis:1985:SAS

[216] Marija J. Norušis. SPSS-X Advanced Statistics Guide. McGraw-Hill, Inc. and SPSS, Inc., New York, NY, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. ISBN 0-07-046548-7 (paperback). v + 505 pp. LCCN HA32 .N6515 1985.

Norusis:1985:SPA

[217] Marija J. Norušis. SPSS/PC+ Advanced Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. ISBN 0-918469-57-0 (manual). LCCN HA 32 N6634 1988. Version 3.1. 6 computer disks, plus reference card (7.5 × 11 in.).

Norusis:1985:SPI

[218] Marija J. Norušis. SPSS/PC+ for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. Program files (IBM PC XT) on 11 disks.

Sola:1985:RSP

[219] Steven Anthony Sola. Review of SPSS/ PC from SPSS, Inc. Dr. Dobb’s Journal of Software Tools, 10(4):84–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Spaulding:1985:ETT

[220] Diane Lynn Spaulding. The effectiveness of two teaching methods in computer terminal skills acquisition with SPSS utilization. Thesis (m.s.), University of Southern Mississippi, Hattiesburg, MS, USA, 1985. iv + 51 pp.

SPSS:1985:SG

[221] SPSS, Inc. SPSS Graphics. SPSS, Inc.

and McGraw-Hill, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and New York, NY, USA, 1985. ISBN 0-07-046554-1 (paperback). iii + 305 pp. LCCN HA32 .S62 1985.

SPSS:1985:SSA

[222] SPSS, Inc. SPSS Statistical Algorithms.

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. ISBN 0-918469-17-1. vi + 225 pp.

SPSS:1985:SUCa

[223] SPSS, Inc. SPSSx user code: adding user programs to SPSSx on Digital VAX/ VMS systems. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. ix + 111 pp.

SPSS:1985:SUCb

[224] SPSS, Inc. SPSS^{X} User Code: Adding User Programs to SPSS^{X} on IBM MVS and CMS systems. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. viii + 201 pp.

SPSS:1985:SXT

[225] SPSS, Inc. SPSS X Tables. McGraw-Hill, Inc. and SPSS, Inc., New York, NY, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1985. ISBN 0-07-046558-4 (paperback). xviii + 200 pp. LCCN HA32 .S66 1985.

Stene:1985:ISU

[226] John W. Stene and Peter R. Del Favero. Introduction to SPSS using the PDP 11-44 and POISE. ??, Yankton, SD, USA, 1985. ii + 61 + [3] pp.

Clogg:1986:LLM

[227] Clifford C. Clogg and Mark P. Becker. Log-linear modeling with SPSS. In Allen [524], page ?? ISBN 0-444-70018-8 (paperback). LCCN QA276.4

.S95 1985. URL http://catalog.
hathitrust.org/api/volumes/oclc/
13455449.html;   http://www.dtic.
mil/dtic/tr/fulltext/u2/a166019.
pdf;   http://zbmath.org/?q=an:
0647.62008.

Nocks:1986:LUS

[228] Elaine C. Nocks and Gilles O. Einstein. Learning to Use the SPSS Batch System. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1986. ISBN 0-13-528019-2 (paperback). xviii + 151 pp. LCCN HA 32 N63 1986.

Norusis:1986:SAS

[229] Marija J. Norušis. SPSS/PC+: Advanced Statistics: For the IBM PC/XT/ AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. various pp. LCCN HA 32 N663 1986.

Norusis:1986:SGDa

[230] Marija J. Norušis. The SPSS Guide to Data Analysis. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-24-4 (paperback). xiii + 402 pp. LCCN HA32 .N65181 1986.

Norusis:1986:SGDb

[231] Marija J. Norušis. The SPSS Guide to Data Analysis: Instructor’s Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. iv + 188 pp.

Norusis:1986:SIP

[232] Marija J. Norušis. SPSS/PC+ for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. 643 pp. LCCN HA32 .N667 1986.

Norusis:1986:SPAa

[233] Marija J. Norušis. SPSS/PC+ Advanced Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. Program files (IBM PC) on 3 computer disks manual.

Norusis:1986:SPAb

[234] Marija J. Norušis. SPSS/PC+: Advanced Statistics for the IBM PC/XT/ AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986.

ISBN 0-918469-15-5 (paperback). various pp. LCCN HA 32 N663 1986.

Norusis:1986:SPI

[235] Marija J. Norušis. SPSS/PC+ for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-14-7 (v.1), 0-918469-40-6 (v.2), 0-918469-41-4 (v.3), 0-918469-36-8, 0-918469-36-8 (v.4), 0-918469-35-X (v.5). 643 pp. LCCN HA32 .N667 1986. Program files on 8 computer disks.

Salavitabar:1986:ACD

[236] Hadi Salavitabar. Accessing COMPUS-TAT data through SPSS. Technical report, State University of New York College at New Paltz, New Paltz, NY, USA, 1986. 8 + 4 pp.

SPSS:1986:SDF

[237] SPSS, Inc. SPSS-X Data File Formats: IBM 370 Version. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. iv + 28 + [1] pp. LCCN HA 32 S655 1986.

SPSS:1986:SG

[238] SPSS, Inc. SPSS Graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, revised edition, 1986. ISBN 0-918469-23-6. iii + 343 pp. LCCN HA 32 S62 1986.

SPSS:1986:SGR

[239] SPSS, Inc. SPSS Graphics, Release 1.1 Update. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-26-0. ii + 36 pp. LCCN HA32 .S62 1985 Suppl.

SPSS:1986:SP

[240] SPSS, Inc. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. Program files (IBM PC) on 2 computer disks, plus update booklet.

SPSS:1986:SPGa

[241] SPSS, Inc. SPSS/PC+ Graphics for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. 64 pp.

SPSS:1986:SPGb

[242] SPSS, Inc. SPSS/PC+ Graphics V2.0:

for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-27-9. 64 pp. LCCN ??

SPSS:1986:SPGc

[243] SPSS, Inc. SPSS/PC+ Graph-in-the-box. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-945830-00-9 (Handbook) 0-918469-49-X (Manual). LCCN ?? Version 3.1. 2 computer disks + 1 owner’s handbook (182 p.) + 1 command summary card (15x23 cm).

SPSS:1986:SPGd

[244] SPSS, Inc. SPSS/PC+ Graphics Featuring Harvard Graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-923967-01-X (V3.1 manual). LCCN ?? Version 3.0. 7 computer disks + 1 Harvard graphics manual (429 p.) + 1 autographix slide service manual (31 p.) + 1 autographix slide service manual (31 p.) + 1 Version 2.13 release notes manual (61 p.) + SPSS/PC+ graphics 3.1 manual (64 p.).

SPSS:1986:SPIa

[245] SPSS, Inc. SPSS/PC + for the IBM PC/XT/AT: Data Entry. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-35-X. various pp. LCCN ??

SPSS:1986:SPIb

[246] SPSS, Inc. SPSS/PC + for the IBM PC/XT/AT: Entering Data using SPSS/PC + DE. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. [2] + 12 pp.

SPSS:1986:SPT

[247] SPSS, Inc. SPSS/PCp+s tables. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-16-3. 643 pp. LCCN HA32 .N667 1986. US$14.95.

SPSS:1986:SSA

[248] SPSS, Inc. SPSS Statistical Algorithms. SPSS Inc., reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, revised edition, 1986. vi + 225 pp.

SPSS:1986:SSG

[249] SPSS, Inc. Supplement to SPSS graphics. Number 8 in SPSS Inc. report. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-26-0. 36 pp. LCCN HA32 .S62 1985 Suppl.

SPSS:1986:STI

[250] SPSS, Inc. SPSS/PC+ Tables for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-16-3. 210 pp. LCCN HA32 .N667 1986.

SPSS:1986:SXC

[251] SPSS, Inc. SPSS X Capture: Accessing VAX DATATRIEVE in SPSS X. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-30-9. 100 pp. LCCN ??

SPSS:1986:SXU

[252] SPSS, Inc. SPSS X User’s Guide.

McGraw-Hill, Inc. and SPSS, Inc., New York, NY, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, second edition, 1986. ISBN 0-07-046553-3 (MH), 0-918469-18-X (SPSS). vii + 988 pp. LCCN HA32 .S67 1986 Ref.

SPSS:1986:TSP

[253] SPSS, Inc. TRANSLATE SPSS/PC+ for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1986. ISBN 0-918469-34-1. [5] + 20 pp. LCCN ??

Wilke:1986:ANS

[254] Helmut Wilke. Eine Analyse der Nutzung statistischer Programmpakete in der Sozialforschung [English: An Analysis of the Utilization of Statistical Program Packages in Social Research]. Quorum, Berlin, Germany, 1986. ISBN 3-88726-016-3. 267 pp. LCCN HA32 .W54 1986.

Comtrain:1987:SPF

[255] Comtrain, Inc. and Vasco Corporation.

SPSS/PC+ Fast Start. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-65-1 (manual). LCCN H 61.3 S6 1988. 2 computer disks.

Einstein:1987:LUS

[256] Gilles O. Einstein and Elaine C. Nocks. Learning to Use SPSS^{x}. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1987. ISBN 0-13-528050-8. viii + 183 pp. LCCN HA32 .E451 1987.

Frude:1987:GSPa

[257] Neil Frude. A guide to SPSS/PC+. Springer-Verlag and Macmillan Education, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Basingstoke, Hampshire, England, 1987. ISBN 0-387-91312-2. xv + 279 pp. LCCN HA 32 F78 1987.

Frude:1987:GSPb

[258] Neil Frude. A guide to SPSS/ PC+. Macmillan Education, Basingstoke, Hampshire, England, 1987. ISBN 0-333-44717-4 (paperback). 320 pp. LCCN HA32.F78 1987.

Hedderson:1987:SMS

[259] John Hedderson. SPSS-X Made Simple. Wadsworth Publishing Company, Belmont, CA, USA, 1987. ISBN 0-534-07458-8 (paperback). xvi + 253 pp. LCCN HA32 .H431 1987.

Norusis:1987:SGD

[260] Marija J. Norušis. The SPSS Guide to Data Analysis for SPSS with Additional Instructions for SPSS/PC. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-42-2 (paperback). xiii + 428 pp. LCCN HA32 .N65191 1987 Desk.

Norusis:1987:SPa

[261] Marija J. Norušis. SPSS/PC+. New England Software, Greenwich, CT, USA, 1987. Version 2.0. 8 computer disks.

Norusis:1987:SPb

[262] Marija J. Norušis. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-55-4 (Base manual), 0-918469-61-9 (Installation guide), 0-918469-58-9 (Getting started guide), 0-918469-57-0 (Advanced statistics guide), 0-918469-43-0 (Data entry II guide), 0-918469-47-3 (statistics guide), 0-918469-43-0 (Data entry II guide), 0-918469-47-3 (Data entry II Entering data guide), 0-918469-46-5 (Data entry II Getting started guide), 0-918469-65-1 (Fast start guide), 0-918469-27-9 (Graphics manual), 0-918469-39-2 (Mapping guide), 0-918469-56-2 (Tables guide), 0-918469-60-0 (Tables Getting started guide), 0-918469-44-9 (Trends guide), 0-918469-59-7 (Trends Getting started guide). LCCN HA32 .N6674 1988. Version 2.0. 24 computer disks + 1 base manual + 1 installation guide + 1 getting started guide + 1 advanced statistics guide + 3 data entry guides + 1 fast start guide + 1 graphics manual + 2 3 data entry guides + 1 fast start guide + 1 graphics manual + 2 tables guides + 2 trends guides + 1 mapping guide.

SPSS:1987:SDE

[263] SPSS, Inc. SPSS Data Entry II. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-923967-19-2. LCCN QA76.9.D3. various.

SPSS:1987:SMI

[264] SPSS, Inc. SPSS/PC+ Mapping: For the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-39-2. 40 pp. LCCN HA 32 S65 1987.

SPSS:1987:SPD

[265] SPSS, Inc. SPSS/PC+ Demo Guided Tour. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. 2 computer disks sales pamphlet.

SPSS:1987:SPG

[266] SPSS, Inc. SPSS/PC+ Graph-in-the-box. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. 26 pp. Version 3.0.

SPSS:1987:SPI

[267] SPSS, Inc. SPSS/PC + for the IBM PC/XT/AT: update. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. [6] + 25 + [1] pp.

SPSS:1987:SPT

[268] SPSS, Inc. SPSS/PC+ tables. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-56-2. LCCN HA 32 S626 1988. Version 2.0. 2 computer disks.

SPSS:1987:STI

[269] SPSS, Inc. SPSS/PC+ Trends: For the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. ISBN 0-918469-44-9 (paperback). 299 pp. LCCN HA 32 S627 1987.

SPSS:1987:STP

[270] SPSS, Inc. SPSS-X Tables primer. SPSS data analysis software. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1987. i + 30 pp.

Weiss:1987:SSA

[271] Neil A. Weiss. SPSS Supplement to Accompany Weiss/Hassett Introductory statistics, second edition. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-09587-4 (paperback). x + 198 + 48 + 6 pp. LCCN QA276.12 .W45 1987 Suppl.

Anonymous:1988:SAS

[272] Anonymous. SPSS announces SPSS-X for Cray Unicos. Computational Statistics & Data Analysis, 7(1):100– 101, August 1988. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900205.

ComTrain:1988:SFS

[273] ComTrain, Inc. SPSS/PC+ Fast Start.

SPSS Europe B.V., Gorinchem, The Netherlands, 1988. ISBN 0-918469-65-1. 34 pp. LCCN H 61.3 S6 1988.

Joreskog:1988:LGP

[274] K. G. Joreskog and Dag Sorbom. LIS-REL 7: A Guide to the Program and Applications. Prentice-Hall, Inc. and SPSS, Inc., Englewood Cliffs, NJ, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-13-177965-6 (P-H), 0-918469-82-1 (SPSS). xx + 342 pp.

Kahler:1988:SAE

[275] Wolf-Michael Kahler. SPSS für Anfanger: eine Einführung in das Datenanalysesystem [English: SPSS for Beginners: An Introduction to the Data Analysis System]. Friedrich Vieweg und Sohn, Braunschweig, Germany, second revised edition, 1988. 272 pp.

Kocklauner:1988:ARS

[276] G. (Gerhard) Kockläuner. Angewandte Regressionsanalyse mit SPSS [English: Applied Regression Analysis with SPSS]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1988. ISBN 3-528-04605-8. viii + 170 pp. LCCN QA278.2 .K63 1988.

Norusis:1988:SGDa

[277] Marija J. Norušis. The SPSS Guide to Data Analysis for SPSSx. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-42-2. xiii + 402 pp. LCCN HA32 .N65191 1987.

Norusis:1988:SGDb

[278] Marija J. Norušis. The SPSS Guide to Data Analysis for SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-33-3. xiv + 434 pp. LCCN HA32 .N651891 1988.

Norusis:1988:SIS

[279] Marija J. Norušis. SPSS-X Introductory Statistics Guide for SPSS-X Release 3. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-54-6 (paperback). xiii + 370 pp. LCCN HA 32 N68 1988.

Norusis:1988:SP

[280] Marija J. Norušis. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. Version 2.0. 8 computer disks.

Norusis:1988:SPA

[281] Marija J. Norušis. SPSS/PC+ Advanced Statistics V2.0 for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N.

Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-57-0 (paperback). various pp. LCCN HA 32 N6634 1988.

Norusis:1988:SPI

[282] Marija J. Norušis. SPSS/PC+ for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-46-5 (Data Entry II Getting started guide) 0-918469-68-6. LCCN (Installation guide) 0-918469-67-8 (Update manual). Version 3.0. 18 computer disks + 1 update manual + 1 installation guide + 1 Data Entry II Getting started guide + 1 blank disk. guide + 1 Data Entry II Getting started guide + 1 blank disk.

Norusis:1988:SPSa

[283] Marija J. Norušis. SPSS/PC+ studentware. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. xiii + 61 pp. Preview version.

Norusis:1988:SPSb

[284] Marija J. Norušis. SPSS/PC+ Studentware. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. Complimentary version. 4 computer disks and fact sheet.

Norusis:1988:SPSc

[285] Marija J. Norušis. SPSS-PC+ Studentware. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-73-2. xiv + 434 pp. LCCN HA32 N6674 1988.

Norusis:1988:SPV

[286] Marija J. Norušis. SPSS/PC+ V2.0 Base Manual: For the IBM PC/XT/ AT and PS/2. SPSS. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-55-4 (paperback). 606 pp. LCCN HA32 .N6674 1988.

SPSS:1988:EDS

[287] SPSS, Inc. Entering Data with SPSS Data Entry II: For the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-923967-22-2. 12 pp. LCCN ??

SPSS:1988:GSSa

[288] SPSS, Inc. Getting started with SPSS/ PC+ for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-923967-17-6. 10 pp. LCCN ??

SPSS:1988:GSSb

[289] SPSS, Inc. Getting started with SPSS/ PC+ Trends for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-923967-18-4. 11 pp. LCCN ??

SPSS:1988:SDE

[290] SPSS, Inc. SPSS Data Entry II for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-47-3 (Entering Data), 0-918469-43-0 (SPSS Data Entry II), 0-918469-46-5 (Getting Started). various pp. LCCN H 62.3 S181 1987, H 61.3 S18 1987.

SPSS:1988:SDF

[291] SPSS, Inc. SPSS-X data file formats: Digital VAX/VMS version. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. 33 pp.

SPSS:1988:SP

[292] SPSS, Inc. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-71-6 (Installation guide), 0-918469-83-X (Getting started manual), 0-918469-91-0 (Update manual), 0-918469-55-4 (Base manual). Version 3.1. 10 computer disks + 1 Getting started manual (10 p.) + 1 Update manual ([156] p.) with reference card SPSS/PC+ 3.11 manual ([156] p.) with reference card SPSS/PC+ 3.11 folded to 7.5 x 11 in. + 1 Base manual ([607] p.) with reference card SPSS/PC+ 2.0 folded to 7.5 × 10.5 in.

SPSS:1988:SPA

[293] SPSS, Inc. SPSS/PC+ advanced statistics V2.0: for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. various pp.

SPSS:1988:SPC

[294] SPSS, Inc. SPSS/PC+ [computer file].

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-83-X (Getting Started), 0-918469-67-8 (V3.0 Update Manual), 0-918469-68-6 (V3.0 Installation Guide). LCCN ?? Version 3.0. Seven computer disks, plus Getting Started Guide (10 pp.), plus Update Manual.

SPSS:1988:SPF

[295] SPSS, Inc. SPSS/PC+ Fast Start for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-923967-24-9. 34 pp. LCCN ??

SPSS:1988:SPTa

[296] SPSS, Inc. SPSS/PC+ Tables [computer file]. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-56-2. LCCN HA 32 S626 1988.

SPSS:1988:SPTb

[297] SPSS, Inc. SPSS/PC+ Trends [computer file]. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. Two computer disks.

SPSS:1988:SPV

[298] SPSS, Inc. SPSS/PC+ V2.0 base manual: for the IBM PC/XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. various pp.

SPSS:1988:STI

[299] SPSS, Inc. SPSS Trends II. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ISBN 0-918469-72-4. 512 pp. LCCN ?? US$29.95.

SPSS:1988:STX

[300] SPSS, Inc. SPSS-X Trends. SPSS, Inc. and Prentice-Hall, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and Englewood Cliffs, NJ, USA, 1988. ISBN 0-918469-52-X (SPSS), 0-13-177825-0 (PH). 343 pp. LCCN HA 30.3 S67 1988. US$22.00.

SPSS:1988:SUG

[301] SPSS, Inc. SPSS-X User’s Guide. SPSS Europe B.V., Gorinchem, The Netherlands, third edition, 1988. ISBN 0-918469-51-1 (paperback). 1072 pp. LCCN HA33 .S17 1988.

SPSS:1988:SXC

[302] SPSS, Inc. SPSS X capture: for DATA-TRIEVE. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. 106 pp.

SPSS:1988:SXT

[303] SPSS, Inc. SPSS-X Tables. SPSS, Inc.

and Prentice-Hall, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and Englewood Cliffs, NJ, USA, second edition, 1988. ISBN 0-918469-53-8 (SPSS), 0-13-177833-1 (PH). 214 pp. LCCN HA 32 S66 1988. US$22.00.

SPSS:1988:ULU

[304] SPSS, Inc. USERPROC LISREL: using LISREL VI within SPSS-X. SPSS Inc. reports. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1988. ii + 37 pp.

Wong:1988:MIS

[305] Kwok-Leung Ken Wong. Man-machine interface for SPSS development [microform]. Thesis (m.sc.), University of Calgary, Calgary, Alberta, Canada, 1988. 2 microfiches.

Anonymous:1989:SAI

[306] Anonymous. SPSS announces interfaces to Informix and Sybase. Computational Statistics & Data Analysis, 8(2):211, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900108.

Anonymous:1989:SII

[307] Anonymous. SPSS Inc. introduces SPSS Graphics with databases interfaces for Informix and Sybase. Computational Statistics & Data Analysis, 8(2):212, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738990011X.

Anonymous:1989:SISa

[308] Anonymous. SPSS introduces SPSS/PC + v3.0. Computational Statistics & Data Analysis, 7(3):318– 319, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900340.

Anonymous:1989:SIV

[309] Anonymous. SPSS increases value of SPSS/PC + by bundling its graphin-the-box option. Computational Statistics & Data Analysis, 8(2): 212–213, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900121.

Backhaus:1989:EMS

[310] Klaus Backhaus. Entwicklung einer Marketing-Konzeption mit SPSS/PC+ [English: Development of a Marketing Idea with SPSS/PC+]. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-51019-2 (New York). xvii + 208 pp. LCCN HF5415.125.B33 1989.

Joreskog:1989:LGP

[311] K. G. Joreskog and Dag Sorbom. LIS-REL 7: A Guide to the Program and Applications. SPSS, Inc., Suite 3300, 444 N.

Michigan Ave., Chicago, IL, USA, second edition, 1989. ISBN 0-918469-94-5. xx + 342 pp. US$30.00.

Keyes:1989:SAG

[312] Barry Keyes. Statistical analysis and graphics SPSS: decision support application review. Meetings Internationale, Ltd., Louisville, KY, USA, 1989. 1 sound cassette.

Keyes:1989:SDA

[313] Barry Keyes. SPSS data analysis/ graphics update. Meetings Internationale, Ltd., Louisville, KY, USA, 1989. 1 sound cassette.

Kolles:1989:SAM

[314] Harry Kolles. Statistisches Auswertungen in der Medizin: ein praktischer Ratgeber für Doktoranden und Wissenschaftler mit zahlreichen Beispielen und einer Einführung in die computergestützte Auswertung mit dem Statistikprogramm SPSS [English: Statistical Evaluations in Medicine: a Practical Advisor for Medical Doctors and Scientists with Numerous Examples and an Introduction in Computer-Supported Evaluation with the Statistics Program SPSS]. Jungjohann, Neckarsulm, Germany, 1989. ISBN 3-8243-1041-4. 120 pp. LCCN ??

Magidson:1989:SPC

[315] Jay Magidson. SPSS/PC+ CHAID.

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-923967-04-4. 43 pp. LCCN ??

Norusis:1989:SP

[316] Marija J. Norušis. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-57-0 (Advanced statistics manual) 0-918469-67-8 (Update manual) 0-918469-55-4 (Base manual). Version 3.0. 17 computer disks pagings: + 1 base manual + 1 Advanced statistics manual + 1 Advanced statistics manual + installation instructions.

SPSS:1989:GSSa

[317] SPSS, Inc. Getting Started with SPSS-X on VAX/VMS. SPSS International, Gorinchem, The Netherlands, 1989. ISBN 0-918469-80-5. 261 pp. LCCN HA 32 G48 1989.

SPSS:1989:GSSb

[318] SPSS, Inc. Getting Started with SPSS Data Entry II: For the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, revised edition, 1989. ISBN 0-923967-21-4. 11 pp. LCCN ??

SPSS:1989:SDE

[319] SPSS, Inc. SPSS-X Data Entry for VAX/VMS. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-84-8. 153 pp. LCCN QA 76.76 O63 S62 1989.

SPSS:1989:SGC

[320] SPSS, Inc. SPSS Graphics Coordinator’s Notebook. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. various pp.

SPSS:1989:SGRa

[321] SPSS, Inc. SPSS Graphics Rerun Facility: VAX/VMS Version. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. various pp.

SPSS:1989:SGRb

[322] SPSS, Inc. SPSS Graphics: Release 2.0 Update. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. iv + 24 pp.

SPSS:1989:SM

[323] SPSS, Inc. SPSS for the Macintosh.

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. Version 4.0.5. 10 computer disks.

SPSS:1989:SP

[324] SPSS, Inc. SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-91-0 (Update Manual) 0-918469-71-6 (Installation Guide), 0-918469-83-X (Getting Started Manual) 0-918469-57-0 (Advanced Statistics Manual) 0-918469-55-4 (Base Manual). Version 3.1. 16 computer disks programs + 1 update manual + 1 base manual + 1 advanced statistics manual + 1 installation guide (11 p.) + 1 getting started guide (10 p.).

SPSS:1989:SPA

[325] SPSS, Inc. SPSS/PC+ Advanced Statistics [computer file]. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. Three computer disks.

SPSS:1989:SPC

[326] SPSS, Inc. SPSS/PC [computer file].

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-91-0. LCCN HA 32 S25 1989. Version 3.1. Three computer disks.

SPSS:1989:SPG

[327] SPSS, Inc. SPSS/PC+ Graphics V3.0:

featuring Harvard Graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-95-3. viii + 56 pp. LCCN ??

SPSS:1989:SPI

[328] SPSS, Inc. SPSS/PC+ for the IBM PC/ XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-71-6 (Installation guide) 0-918469-91-0 (Update manual). LCCN 0-918469-43-0 (Data Entry II guide) 0-923967-05-2 (Data Entry II Getting started guide) 0-918469-47-3 (Data Entry II Entering data Getting started guide) 0-918469-47-3 (Data Entry II Entering data guide) 0-923967-01-X (Graphics manual). Version 3.1. 9 computer disks guide + 3 data entry guides + 2 templates + 1 graphics manuals. guide + 3 data entry guides + 2 templates + 1 graphics manuals.

SPSS:1989:SPU

[329] SPSS, Inc. SPSS/PC+ update for V3.0 and V3.1 for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1989. ISBN 0-918469-91-0. various pp. LCCN HA 32 S25 1989.

WSDSHS:1989:SMS

[330] Washington State, Dept. of Social and Health Services, Olympia, WA, USA. SPSS manual: Social Service Payment System, 1989. various pp.

Anonymous:1990:SAM

[331] Anonymous. SPSS announces Macintosh package. Computational Statistics & Data Analysis, 9(1):161, January 1990. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/016794739090078V.

Bryman:1990:QDA

[332] Alan Bryman and Duncan Cramer.

Quantitative Data Analysis for Social Scientists. Routledge, London, UK; New York, NY, USA, 1990. ISBN 0-415-02664-4 (hardcover), 0-415-02665-2 (paperback). xiv + 290 pp. LCCN HA32 .B79 1990.

Lee:1990:SPG

[333] Ivy Lee, Minako K. Maykovich, and Richard J. Larsen. SPSS/PC+ Guide, Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1990. ISBN 0-13-844234-7. 184 pp. LCCN HA32.L44 1990.

Norusis:1990:SASa

[334] Marija J. Norušis. SPSS Advanced Statistics User’s Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-90-2. 285 pp. LCCN HA32 .N6514 1990.

Norusis:1990:SASb

[335] Marija J. Norušis. SPSS Advanced Statistics: Student Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-03-6 (paperback). 506 pp. LCCN HA32 .N65155 1990.

Norusis:1990:SBS

[336] Marija J. Norušis. SPSS Base System User’s Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-63-5. 520 pp. LCCN HA32 .N658 1990 Annex Controlled Circulation.

Norusis:1990:SGD

[337] Marija J. Norušis. The SPSS Guide to Data Analysis for Release 4. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-08-7. 470 pp. LCCN HA32 .N65186 1990. Revision B, revised and updated.

Norusis:1990:SIS

[338] Marija J. Norušis. SPSS Introductory Statistics: Student Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-02-8. 420 pp. LCCN HA32 .N6758 1990.

Norusis:1990:SPAa

[339] Marija J. Norušis. SPSS/PC+ advanced statistics 4.0 for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-13-3. various pp. LCCN HA32.N662 1990.

Norusis:1990:SPAb

[340] Marija J. Norušis. SPSS/PC+ Advanced Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-13-3 (Manual). LCCN ?? Version 4.0. 5 computer disks reference card (7.5 × 11 in.).

Norusis:1990:SPBa

[341] Marija J. Norušis. SPSS/PC+ 4.0 base manual for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-11-7. various pp. LCCN ??

Norusis:1990:SPBb

[342] Marija J. Norušis. SPSS/PC+ Base system. SPSS, Inc., Suite 3300, 444 N.

Michigan Ave., Chicago, IL, USA, 1990. Version 4.0. 8 computer disks + manual.

Norusis:1990:SPS

[343] Marija J. Norušis. SPSS/PC+ statistics 4.0 for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-12-5. various pp. LCCN HA32.N6672 1990. 3 computer disks, plus reference card (7.5 × 11 in.).

Rowland:1990:BCA

[344] David Rowland, Daniel Arkkelin, and Larry Crisler. Computer-Based Data Analysis: Using SPSSx in the Social and Behavioral Sciences. Nelson-Hall, Chicago, IL, USA, 1990. ISBN 0-8304-1181-X. ix + 213 pp. LCCN HA32.R68 1991.

SPSS:1990:GSS

[345] SPSS, Inc. Getting started with SPSS/ PC + tables for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-23-0. 11 pp. LCCN ??

SPSS:1990:LPU

[346] SPSS, Inc. LISREL 7 and PRELIS User’s Guide and Reference. SPSS, Inc. and Prentice-Hall, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and Englewood Cliffs, NJ, USA, 1990. ISBN 0-918469-70-8 (SPSS), 0-13-177874-9 (PH). LCCN HA32 .S623 1990. US$28.00.

SPSS:1990:MSP

[347] SPSS, Inc. Map 4.0 SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-15-X. LCCN GA102.4. 1 computer disk.

SPSS:1990:SC

[348] SPSS, Inc. SPSS Categories. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-93-7. various pp. LCCN HA32 .S547 1990.

SPSS:1990:SGRa

[349] SPSS, Inc. SPSS Graphics Release 2.1.3 for SunOS 4.1. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. various pp.

SPSS:1990:SGRb

[350] SPSS, Inc. SPSS Graphics: Release 2 Update. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-99-6. iv + 25 pp. LCCN HA 32 S621 1990.

SPSS:1990:SIC

[351] SPSS, Inc. SPSS for IBM CMS: Operations Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-96-1 (paperback). 138 pp. LCCN HA32.S56 1990.

SPSS:1990:SIM

[352] SPSS, Inc. SPSS for IBM MVS: Operations Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-09-5. 140 pp. LCCN HA32 .S57 1990.

SPSS:1990:SMO

[353] SPSS, Inc. SPSS for the Macintosh: Operations Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-98-8. 159 pp. LCCN ??

SPSS:1990:SOOa

[354] SPSS, Inc. SPSS for OS/2: Operations Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-918469-92-9 (paperback). 196 pp. LCCN ??

SPSS:1990:SOOb

[355] SPSS, Inc. SPSS for OS/2: Operations Guide, Release 4.1. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-27-3. 235 pp. LCCN HA32 .S575 1990.

SPSS:1990:SPG

[356] SPSS, Inc. SPSS/PC+ Graphics for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-25-7. viii + 64 pp. LCCN ??

SPSS:1990:SPI

[357] SPSS, Inc. SPSS/PC+ for the IBM PC/ XT/AT. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. Version 4.0. 3 computer disks.

SPSS:1990:SPP

[358] SPSS, Inc. SPSS PC Plus Base Manual 4.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1990. ISBN 0-13-178112-X. ?? pp. LCCN ?? US$22.00.

SPSS:1990:SPTa

[359] SPSS, Inc. SPSS/PC+ Tables V2.0: for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-20-6 (paperback). various pp. LCCN ??

SPSS:1990:SPTb

[360] SPSS, Inc. SPSS/PC+ Trends for the IBM PC/XT/AT and PS/2. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-923967-14-1. various pp. LCCN HA32.S628 1990.

SPSS:1990:SRG

[361] SPSS, Inc. SPSS Reference Guide. Prentice-Hall, Inc. and SPSS, Inc., Englewood Cliffs, NJ, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1990. ISBN 0-13-177858-7 (PH), 0-918469-62-7 (SPSS). 949 pp. LCCN HA32 .S634 1990. US$41.00.

SPSS:1990:SUO

[362] SPSS, Inc. SPSS for UNIX: Operations Guide. SPSS, Inc. and Prentice-Hall, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and Englewood Cliffs, NJ, USA, 1990. ISBN 0-918469-97-X (SPSS), 0-13-177981-8 (PH). 152 pp. LCCN HA32 .S625 1990 Reserve Controlled Circulation. US$12.00.

SPSS:1990:SVV

[363] SPSS, Inc. SPSS for VAX/VMS: Operations Guide. SPSS International, Gorinchem, The Netherlands, 1990. ISBN 0-923967-07-9. 258 pp. LCCN HA32 .S59 1990.

Friesen:1991:CGS

[364] Blake Friesen. A Concise Guide to Statistical Software Packages: An Introduction to Minitab, SPSS/PC+, SAS, NWA Statpak, and Lotus 1-2-3. Harcourt Brace Jovanovich Canada, Toronto, Ontario, Canada, 1991. ISBN 0-03-997801-X. 87 pp. LCCN ??

Hedderson:1991:SMS

[365] John Hedderson. SPSS/PC+ Made Simple. Wadsworth Publishing Company, Belmont, CA, USA, 1991. ISBN 0-534-14376-8, 0-534-98285-9. xvi + 271 pp. LCCN HA32 .H42 1991 DESK.

Levine:1991:GSA

[366] Gustav Levine. A Guide to SPSS for Analysis of Variance. Lawrence Erlbaum Associates, Hillsdale, NJ, USA, 1991. ISBN 0-8058-0939-2, 0-8058-0941-4 (paperback). xvii + 151 pp. LCCN HA31.35 .L48 1991.

Norusis:1991:SGD

[367] Marija J. Norušis. The SPSS Guide to Data Analysis for SPSS/PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, second edition, 1991. ISBN 0-923967-34-6. 499 pp. LCCN HA32 .N65188 1991.

Norusis:1991:SPSa

[368] Marija J. Norušis. SPSS/PC+ Studentware Plus. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. Version 1.0. 5 computer disks.

Norusis:1991:SPSb

[369] Marija J. Norušis. SPSS/PCb+s Studentware Plus for Business. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. ISBN 0-923967-33-8. 604 pp. LCCN HF5548.2. 10 computer disks and reference card.

Norusis:1991:SPSc

[370] Marija J. Norušis. SPSS/PCb+s Studentware Plus. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. 483 pp.

Norusis:1991:SPSd

[371] Marija J. Norušis. SPSS/PC+ Studentware Plus: Instructor’s Manual. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. 166 pp.

Norusis:1991:SPSe

[372] Marija J. Norušis. SPSS/PC+ Studentware Plus. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. ISBN 0-923967-32-X (text). LCCN HA32. Version 1.0. 10 computer disks.

Norusis:1991:SPSf

[373] Marija J. Norušis. SPSS/PC+ Studentware for Business. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. ISBN 0-923967-30-3, 0-923967-31-1. 604 pp. LCCN ??

Norusis:1991:SPSg

[374] Marija J. Norušis. SPSS/PC+ Studentware Plus. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, revised edition, 1991. ISBN 0-923967-28-1, 0-923967-29-X. 483 pp. LCCN ??

Schubo:1991:SKV

[375] Werner Schubo. SPSS Kompakt für die Versionen 3 und 4 [English: SPSS Compact Guide for Versions 3 and 4]. G. Fischer, Stuttgart, Germany, 1991. ISBN 3-437-40257-9. 126 pp. LCCN ??

SPSS:1991:SG

[376] SPSS, Inc. SPSS Graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, revised edition, 1991. ISBN 0-923967-35-4. iii + 343 + 25 pp. LCCN ??

SPSS:1991:SPC

[377] SPSS, Inc. SPSS/PC+ Codebook. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1991. ISBN 0-923967-38-9 (paperback). various pp. LCCN HA32.S624.

SPSS:1991:SSA

[378] SPSS, Inc. SPSS, Statistical Algorithms.

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, second edition, 1991. ISBN 0-918469-89-9. 290 pp. LCCN CM.1.1.05. US$28.00.

West:1991:CPS

[379] Robert West. Computing for Psychologists: Statistical Analysis Using SPSS and MINITAB. Harwood Academic Publishers, Chur, Switzerland; New York, NY, USA, 1991. ISBN 3-7186-5085-1 (hard), 3-7186-5086-X (soft). ix + 234 pp. LCCN BF39 .W45 1991.

Abouserie:1992:SME

[380] R. (Reda) Abouserie. Statistical Methods for Educational and Psychological Research: Basic Concepts and Guide to Data Analysis Using SPSS/PC+ 4.01. School of Education, University of Wales, College of Cardiff, Cardiff, Wales, 1992. ISBN 1-872330-01-0. 125 pp. LCCN HA29.A29 1992.

Brodeur:1992:IIP

[381] Patrice Brodeur. IFT 1990, informatique pour sciences sociales: SPSS/PC+ studentware: notes et références. Technical report, Librairie de l’Université de Montréal, Montréal, Québec, Canada, 1992. ii + 43 pp.

Dometrius:1992:SSU

[382] Nelson C. Dometrius. Social Statistics using SPSS. HarperCollins College Publishers, New York, NY, USA, 1992. ISBN 0-06-041714-5. xx + 540 pp. LCCN HA32 .D66 1992.

Foster:1992:SSP

[383] Jeremy J. Foster. Starting SPSS/PC+: A Beginner’s Guide to Data Analysis. Sigma Press, Wilmslow, England, 1992. ISBN 1-85058-263-7 (paperback). x + 200 pp. LCCN ??

Holzer:1992:MCP

[384] Christian Holzer and Manfred Precht. Multiple comparison procedures for normally distributed ANOVA models in SAS, SPSS, BMDP, and MINITAB. Computational Statistics & Data Analysis, 13(3):351–358, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947392901412.

Kinnear:1992:SMS

[385] Paul R. Kinnear and Colin D. Gray. SPSS/PC+ Made Simple. Lawrence Erlbaum Associates, Hillsdale, NJ, USA, 1992. ISBN 0-86377-297-8 (paperback). viii + 210 pp. LCCN HA32 .K56 1992.

Kosfeld:1992:BRB

[386] Reinhold Kosfeld. Book review: SPSS/PC+ Basics and graphic: Brosius, G. (1988) Einführung und praktische Beispiele. Hamburg: McGraw-Hill Book Company. 582 pp, ISBN 3-89028-132-X, ca. DM 68,-. Computational Statistics & Data Analysis, 13(4):494–495, May 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739290126Z.

Magidson:1992:SPC

[387] Jay Magidson. SPSS/PC+ CHAID, Version 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-84-2. xii + 149 pp. LCCN ??

Norusis:1992:SPA

[388] Marija J. Norušis. SPSS/PCp+s Advanced Statistics Version 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-68-0. xiii + 481 pp. LCCN HA32.N6636 1992.

Norusis:1992:SPB

[389] Marija J. Norušis. SPSS/PC+ Base System User’s Guide, Version 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-66-4. xvii + 910 pp. LCCN HA32.N6637 1992.

Norusis:1992:SPP

[390] Marija J. Norušis. SPSS/PC+ Professional Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-67-2. xi + 236 pp. LCCN ?? Version 5.0.

Norusis:1992:SWA

[391] Marija J. Norušis. SPSS for Windows:

Advanced Statistics Release 5. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-56-7 (paperback). 580 pp. LCCN ??

Norusis:1992:SWB

[392] Marija J. Norušis. SPSS for Windows:

Base System User’s Guide, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-49-4 (paperback). 672 pp. LCCN HA32.N6517 1992.

Norusis:1992:SWP

[393] Marija J. Norušis. SPSS for Windows: Professional Statistics, Release 5. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-50-8 (paperback). 348 pp. LCCN HA32.N65173 1993.

Shin:1992:SPG

[394] Kilman Shin. SPSS/PCb+s Guide: For Use with Applied Statistics for Business and Economics, Allen Webster. Irwin, Homewood, IL, USA, 1992. ISBN 0-256-10959-1. xiii + 389 pp. LCCN HF1017.S5234 1992.

SPSS:1992:GSS

[395] SPSS, Inc. Getting started with SPSS/ PC+. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-71-0. iii + 16 pp. LCCN ?? Version 5.0.

SPSS:1992:ISP

[396] SPSS, Inc. Installing SPSS/PC+ and SPSS Data Entry II. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-72-9. iii + 18 pp. LCCN ?? Version 5.0.

SPSS:1992:SBSa

[397] SPSS, Inc. SPSS/PC+ Base System User’s Guide, Version 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1992. ISBN 0-13-177692-4. US$33.00.

SPSS:1992:SBSb

[398] SPSS, Inc. SPSS Base System Syntax Reference Guide, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-53-2. 792 pp. LCCN HA32 .S545 1992.

SPSS:1992:SGTa

[399] SPSS, Inc. SPSS/PC: Graphics from TriMetrix. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, second edition, 1992. ISBN 0-13-108879-3. ?? pp.

SPSS:1992:SGTb

[400] SPSS, Inc. SPSS/PC+ Graphics from TriMetrix. Prentice-Hall, Inc. and SPSS, Inc., Englewood Cliffs, NJ, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-13-177759-9 (PH), 0-923967-79-6 (SPSS). xviii + 575 pp. US$24.00.

SPSS:1992:SPG

[401] SPSS, Inc. SPSS/PC+ graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. Version 4.0.1. 11 computer disks.

SPSS:1992:SSP

[402] SPSS, Inc. SPS S/PC+ base. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. Version 5.0.2 upgrade. 4 computer disks.

SPSS:1992:SWB

[403] SPSS, Inc. SPSS for Windows: Base System User’s Guide, Release 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1992. ISBN 0-13-178526-5 (paperback). ?? pp. LCCN ??

SPSS:1992:SWH

[404] SPSS, Inc. SPSS for Windows: Help for SPSS/PC+ Users. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-55-9. 14 pp. LCCN HA32.S595 1992.

SPSS:1992:USP

[405] SPSS, Inc. Using SPSS/PC+ Graphics with Harvard Graphics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1992. ISBN 0-923967-80-X (paperback). vii + 24 pp. LCCN ??

Anonymous:1993:BRBg

[406] Anonymous. Book review: Datenanalyse mit SPSS: Wittenberg, Reinhard and Cramer, Hans (1992) Stuttgart: Gustav Fischer Verlag, pp. 200, ISBN 3-8252-1602-0, DM 24.80. Computational Statistics & Data Analysis, 16(2):248, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794739390124C.

Babbie:1993:ASR

[407] Earl R. Babbie. Adventures in Social Research: Data Analysis Using SPSS. The Pine Forge Press social science library. Pine Forge Press, Thousand Oaks, CA, USA, 1993. ISBN 0-8039-9007-3 (paper). xvi + 292 pp. LCCN HA32.B28 1994.

Badarinathi:1993:DPI

[408] Ravija Badarinathi. The Dryden Press Introduction to SPSS. Dryden Press, Fort Worth, TX, USA, 1993. ISBN 0-03-096964-6. 97 pp. LCCN ??

Brodeur:1993:IIP

[409] Patrice Brodeur and Normand Ranger. IFT 1990, informatique pour sciences sociales: SPSS/PC+ studentware: notes et références pour la partie SPSS. Technical report, Librairie de l’Université de Montréal, Montréal, Québec, Canada, 1993. ii + 61 pp.

Fischer:1993:SSI

[410] Henry W. Fischer, III. Social Statistics the IBM-PC, and SPSS/PC+: A Manual for the User. ASA resource materials for teaching. American Sociological Association, Washington, DC, USA, 1993. 300 pp.

Foster:1993:SSP

[411] Jeremy J. Foster. Starting SPSS/ PC+ and SPSS for Windows: A Beginner’s Guide to Data Analysis. Sigma Press, Wilmslow, England, second edition, 1993. ISBN 1-85058-509-1 (paperback). xii + 277 pp. LCCN ??

Frude:1993:GS

[412] Neil Frude. A Guide to SPSS/PC+.

Macmillan Publishing Co., Inc., New York, NY, USA, second edition, 1993. ISBN 0-333-58908-4 (paperback). xvi + 336 pp. LCCN HA32 .F78 1993.

Frude:1993:GSP

[413] Neil Frude. A Guide to SPSS/PC+.

Macmillan Education and Springer-Ver-lag, Basingstoke, Hampshire, England and Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 0-333-58908-4 (Macmillan), 0-387-91442-0 (Springer). xvi + 336 pp. LCCN HA32 .F78 1993.

Hedderson:1993:SMS

[414] John Hedderson and Melinda Fisher.

SPSS Made Simple. Wadsworth Publishing Company, Belmont, CA, USA, second edition, 1993. ISBN 0-534-19992-5. xv + 283 pp. LCCN HA32 .H43 1993.

Huizingh:1993:ISP

[415] Eelko Huizingh. Introduction to SPSS/ PC+5.0 and Data Entry. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52975-0 (paperback). xii + 346 pp. LCCN HA32.H83 1993.

Kim:1993:SCW

[416] Ho jong Kim. Silche charyo wa SPSS/ PCp+s rul hwaryonghan sahoe kwahak tonggye punsok. Samyongsa, Soul Tukpyolsi, Korea, 1-pan edition, 1993. ISBN 89-445-3581-7. 559 pp. LCCN HA29.5.K6K56 1993.

Meyer:1993:SMA

[417] Glenn E. Meyer. SPSS: A Minimalist Approach. Harcourt Brace Jovanovich, Boston, MA, USA, 1993. ISBN 0-03-055392-X. x + 491 pp. LCCN HA32 .M49 1993.

Norusis:1993:SUA

[418] Marija J. Norušis. SPSS for UNIX: Advanced Statistics, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-88-5. xiv + 574 pp. LCCN HA32 .N65166 1993.

Norusis:1993:SUB

[419] Marija J. Norušis. SPSS for UNIX: Base System User’s Guide, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-86-9. xvii + 700 pp. LCCN HA32 .N65167 1993.

Norusis:1993:SUP

[420] Marija J. Norušis. SPSS for UNIX: Professional Statistics, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-87-7. xi + 345 pp. LCCN ??

Norusis:1993:SWA

[421] Marija J. Norušis. SPSS for Windows, Advanced Statistics, Release 6.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-13-178823-X. xiv + 578 pp. LCCN HA32.N65168 1993. US$28.00.

Norusis:1993:SWB

[422] Marija J. Norušis. SPSS for Windows:

Base System User’s Guide 6.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-178856-6. xviii + 828 pp. US$36.00.

Norusis:1993:SWP

[423] Marija J. Norušis. SPSS for Windows: Professional Statistics, Release 6.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-178831-0. xii + 385 pp. US$28.00.

Shurbaji:1993:AAA

[424] Abd al Razzaq Shurbaji. al-Iqtisad al-qiyasi al-tatbiqi: tatbiqiyat ala iqtisadiyat al-duwal al-Arabiyah bi-al-istikhdam Excel, Quatro-Pro, Dbase III Plus and SPSS/PC+. A. Shurbaji, Bayrut, Lubnan, al-tabah 2. edition, 1993. 365 pp.

SPSS:1993:SAS

[425] SPSS, Inc. SPSS for Windows: Advanced Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. Version 6.1. 1 computer disk.

SPSS:1993:SBS

[426] SPSS, Inc. SPSS Base System Syntax Reference Guide, Release 6.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-179391-8. ix + 941 pp. US$36.00.

SPSS:1993:SC

[427] SPSS, Inc. SPSS CHAID. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. Version 6.0.1. 1 computer disk.

SPSS:1993:SGDa

[428] SPSS, Inc. The SPSS Guide to Data Analysis, Release 4.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-178096-4. US$19.00.

SPSS:1993:SGDb

[429] SPSS, Inc. The SPSS Guide to Data Analysis for SPSS/PC+. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, second edition, 1993. ISBN 0-13-178443-9. US$19.00.

SPSS:1993:SLP

[430] SPSS, Inc. SPSS Lisrel 7 and Prelis. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-13-112384-X. xi + 169 pp. LCCN ??

SPSS:1993:SPS

[431] SPSS, Inc. SPSS PC: Students 3.5 and SPSS Guide to Data 4.0 Package. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-125931-8. ?? pp. LCCN ?? US$35.40.

SPSS:1993:SPTa

[432] SPSS, Inc. SPSS/PC+ Tables. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. Version 5.0. 1 computer disk.

SPSS:1993:SPTb

[433] SPSS, Inc. SPSS/PC+ Trends. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-13-178716-0. xii + 465 pp. LCCN ?? US$22.00. Version 5.0. 2 computer disks + 1 Getting started (17 p.).

SPSS:1993:SUA

[434] SPSS, Inc. SPSS for UNIX Advanced Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-107137-8. US$18.00.

SPSS:1993:SUB

[435] SPSS, Inc. SPSS for UNIX Base Systems User’s Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-107145-9. US$24.75.

SPSS:1993:SUP

[436] SPSS, Inc. SPSS for UNIX Professional Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-107152-1. US$18.00.

SPSS:1993:SWD

[437] SPSS, Inc. SPSS 6.0 for Windows: Developer’s Kit. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. 191 pp.

SPSS:1993:SWTa

[438] SPSS, Inc. SPSS for Windows: Tables, Release 5. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-85-0. xii + 259 pp. LCCN HA32.S597 1993.

SPSS:1993:SWTb

[439] SPSS, Inc. SPSS for Windows: Trends, Release 6.0. Prentice-Hall, Inc. and SPSS, Inc., Englewood Cliffs, NJ, USA and Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-13-179383-7 (PH). xii + 353 pp. LCCN HA32.S598 1993. US$22.00.

Babbie:1994:ASR

[440] Earl R. Babbie and Fred Halley. Adventures in Social Research: Data Analysis Using SPSS. The Pine Forge Press Social Science Library. Pine Forge Press, Thousand Oaks, CA, USA, 1994. ISBN 0-8039-9007-3 (paperback). xvi + 292 pp. LCCN HA32 .B28 1994.

Bryman:1994:QDA

[441] Alan Bryman and Duncan Cramer. Quantitative Data Analysis for Social Scientists. Routledge, London, UK; New York, NY, USA, revised edition, 1994. ISBN 0-415-11307-5. ?? pp. LCCN HA32 .B79 1994.

Cramer:1994:ISS

[442] Duncan Cramer. Introducing Statistics for Social Research: Step-by-Step Calculations and Computer Techniques Using SPSS. Routledge, London, UK; New York, NY, USA, 1994. ISBN 0-415-07514-9 (hardback), 0-415-07515-7 (paperback). xiv + 318 pp. LCCN HA32 .C7 1994.

Fan:1994:SSD

[443] Yihua Fan. A systemization of statistical data management/manipulation tasks and a comparison of statistical data management/manipulation capabilities of SAS and SPSS base programs. PhD thesis, University of Virginia, Charlottesville, VA, USA, 1994. xii + 130 pp.

Kinnear:1994:SWM

[444] Paul R. Kinnear and Colin D. Gray. SPSS for Windows Made Simple. LEA, Hove, East Sussex, UK, 1994. ISBN 0-86377-350-8 (paperback). ix + 275 pp. LCCN ??

Norusis:1994:SAS

[445] Marija J. Norušis and SPSS, Inc. SPSS Advanced Statistics 6.1. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-200065-2. xvi + 606 pp.

Norusis:1994:SBS

[446] Marija J. Norušis and SPSS, Inc. SPSS 6.1 Base System User’s Guide, Part 2: UNIX Version. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-438870-4. xiii + 547 pp.

Norusis:1994:SPS

[447] Marija J. Norušis and SPSS, Inc. SPSS Professional Statistics 6.1. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-190125-7. xii + 385 pp. LCCN ??

Norusis:1994:SPW

[448] Marija J. Norušis. SPSS pour Windows:

manuel d’utilisation: système de base, version 6.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. xiv + 330 pp.

Norusis:1994:SSS

[449] Marija J. Norušis. SPSS Professional Statistics 6.1. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-190125-7. xii + 385 pp. LCCN ??

SPSS:1994:SPSa

[450] SPSS, Inc. SPSS/PC+ 5.0 StrateGen.

SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. 141 pp.

SPSS:1994:SPSb

[451] SPSS, Inc. SPSS for Windows: Professional Statistics. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. Version 6.1. 1 computer disk.

SPSS:1994:SSR

[452] SPSS, Inc. SPSS 6.1 Syntax Reference Guide. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-438250-1. ix + 941 pp. LCCN HA32 .S67 1994.

SPSS:1994:ST

[453] SPSS, Inc. SPSS for Windows: Tables. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. Version 6.1. 1 computer disk.

SPSS:1994:SW

[454] SPSS, Inc. SPSS for Windows. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. Version 6.1. 7 computer disks.

SPSS:1995:SBS

[455] SPSS, Inc. SPSS 6.1 Base System User’s Guide, Part 1: UNIX Version. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1995. ISBN 0-13-459561-0. xvii + 341 pp.

Linn:1996:SVE

[456] S. Linn. Sorting variables for each case: a new algorithm to calculate injury severity score (ISS) using SPSS-PC. Statistics in Medicine, 15(7-9):933– 937, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1999:BRS

[457] Nick Longford. Book review: The SPSS guide to the new statistical analysis of data. S. B. Gerber and K. E. Voelkl, Springer-Verlag, New York, 1997. No. of pages: vii + 203. Price: US$32.95. ISBN 0-387-94821-X. Statistics in Medicine, 18(8):1021–1022, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:1999:BRS

[458] Jørgen Holm Petersen. Book review:

SPSS Analysis Without Anguish. S. J. Coakes and L. G. Steed, Wiley, Singapore, 1997. No. of pages: x + 276 + 2 diskettes. Price: £16.99. ISBN 0-471-33869-9. Statistics in Medicine, 18(21):2984–2985, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabroe:1999:BRA

[459] Svend Sabroe. Book review: Adventures in Social Research. Data Analysis Using SPSS for Windows. E. Babbie and F. Halley, Pine Forge Press, 1995. pages: 346 + diskette. Price: £25. ISBN 0-8039-9055-3. Statistics in Medicine, 18(17-18):2550, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

SPSS:19xx:DP

[460] SPSS, Inc. DBMS/COPY Plus.

Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178476-5. US$33.00.

SPSS:19xx:EDS

[461] SPSS, Inc. Entering Data with SPSS Data Entry II. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178377-7. US$6.00.

SPSS:19xx:GSSa

[462] SPSS, Inc. Getting Started With SPSS/ PC+ 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177726-2. US$3.00.

SPSS:19xx:GSSb

[463] SPSS, Inc. Getting Started SPSS/PC+ Tables. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178385-8. US$3.00.

SPSS:19xx:GSSc

[464] SPSS, Inc. Getting Started SPSS Data Entry II. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178211-8. US$3.00.

SPSS:19xx:GSSd

[465] SPSS, Inc. Getting Started with SPSS-X on VAX VMS. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177916-8. US$3.00.

SPSS:19xx:ISV

[466] SPSS, Inc. Installing SPSS/PC+ Version 5.0 and SPSS Data Entry II. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177734-3. US$3.00.

SPSS:19xx:SASa

[467] SPSS, Inc. SPSS Advanced Statistics Student Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178070-0. US$24.00.

SPSS:19xx:SASb

[468] SPSS, Inc. SPSS/PC+ Advanced Statistics 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177718-1. US$24.00.

SPSS:19xx:SASc

[469] SPSS, Inc. SPSS Advanced Statistics User’s Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177940-0. US$22.00.

SPSS:19xx:SBSb

[470] SPSS, Inc. SPSS Base System Syntax Reference Guide, Release 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177668-1. US$36.00.

SPSS:19xx:SBSc

[471] SPSS, Inc. SPSS Base System User’s Guide, Release 4.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177866-8. US$28.00.

SPSS:19xx:SCa

[472] SPSS, Inc. SPSS Categories. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-182320-5. US$33.27.

SPSS:19xx:SCb

[473] SPSS, Inc. SPSS Categories. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177957-5. US$28.00.

SPSS:19xx:SCc

[474] SPSS, Inc. SPSS/PC+ CHAID 5.0.

Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177775-0. US$24.00.

SPSS:19xx:SCd

[475] SPSS, Inc. SPSS/PC+ Codebook.

Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178468-4. US$22.00.

SPSS:19xx:SDE

[476] SPSS, Inc. SPSS Data Entry II. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178195-2. US$22.00.

SPSS:19xx:SG

[477] SPSS, Inc. SPSS Graphics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178450-1. US$36.00.

SPSS:19xx:SIC

[478] SPSS, Inc. SPSS for IBM CMS: Operations Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177973-7. LCCN HA32 .S56 1990. US$12.00.

SPSS:19xx:SIN

[479] SPSS Inc. Newsletter, 19xx. ISSN 0160-7669, 0160-7669. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA.

SPSS:19xx:SIO

[480] SPSS, Inc. SPSS for IBM/MVS: Operations Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178104-9. US$12.00.

SPSS:19xx:SIS

[481] SPSS, Inc. SPSS Introductory Statistics Student Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178062-X. US$22.00.

SPSS:19xx:SLP

[482] SPSS, Inc. SPSS/PC+ LISREL 7 and PRELIS. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177742-4. US$28.00.

SPSS:19xx:SMO

[483] SPSS, Inc. SPSS for Macintosh: Operations Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177999-0. US$12.00.

SPSS:19xx:SOO

[484] SPSS, Inc. SPSS for OS/2: Operations Guide, Release 4.1. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178435-8. US$12.00.

SPSS:19xx:SPS

[485] SPSS, Inc. SPSS/PC+ Professional Statistics 5.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177700-9. US$24.00.

SPSS:19xx:SPT

[486] SPSS, Inc. SPSS/PC+ Tables 4.0.

Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178203-7. US$22.00.

SPSS:19xx:SUA

[487] SPSS, Inc. SPSS for UNIX Advanced Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-107137-8. US$18.00.

SPSS:19xx:SUB

[488] SPSS, Inc. SPSS for UNIX Base Systems User’s Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-107145-9. US$24.75.

SPSS:19xx:SUP

[489] SPSS, Inc. SPSS for UNIX Professional Statistics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-107152-1. US$18.00.

SPSS:19xx:SVO

[490] SPSS, Inc. SPSS for VAX/VMS: Operations Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178088-3. US$12.00.

SPSS:19xx:SWC

[491] SPSS, Inc. SPSS for Windows: CHAID 6.0. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-178849-3. US$24.00.

SPSS:19xx:SWH

[492] SPSS, Inc. SPSS for Windows: Help for SPSS/PC+ Users. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177676-2. US$3.00.

SPSS:19xx:SWT

[493] SPSS, Inc. SPSS for Windows: Tables (for use with Releases 5.0 and 6.1). Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177783-1. US$28.00.

SPSS:19xx:SWU

[494] SPSS, Inc. SPSS 6.1 for Windows: Update. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-182338-8. US$13.27.

SPSS:19xx:SXI

[495] SPSS, Inc. SPSS-X Introductory Statistics Guide to SPSS-X for Release 3. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177841-2. US$19.00.

SPSS:19xx:SXUa

[496] SPSS, Inc. SPSS-X User’s Guide. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, second edition, 19xx. ISBN 0-13-047424-X. US$39.00.

SPSS:19xx:SXUb

[497] SPSS, Inc. SPSS-X User’s Guide.

Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, third edition, 19xx. ISBN 0-13-177817-X. US$39.00.

SPSS:19xx:USG

[498] SPSS, Inc. Using SPSS/PC+ Graphics with Harvard Graphics. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 19xx. ISBN 0-13-177767-X. US$7.00.

UMCC:19xx:ISP

[499] University of Minnesota.Computer Center. Introduction to SPSS, part 1. Technical report, University of Minnesota Computer Center, Minneapolis, MN, USA, 19xx. 1 cassette + 20 min.

Lynch:2003:BRU

[500] Andrew G. Lynch. Book reviews: Using SPSS for Windows. Kristin E. Voelkl and Susan B. Gerber, Springer, New York, 1999. No. of pages: xvi + 228. ISBN 0-387-98563-8. Doing Statistics with SPSS. Alistair W. Kerr, Howard K. Hall and Stephen A. Kozub, Sage Publications, London, 2002. No. of pages: vii + 238. ISBN 0-7169-7385-0. Data Analysis using SPSS for Windows Versions 8-10: a Beginner’s Guide New Edition. Jeremy J. Foster, Sage Publications, London, 2002. No. of pages: xvii + 252. ISBN 0-7169-6927-0. Statistics in Medicine, 22(15):2529–2531, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nielsen:2003:BRS

[501] Tine Nielsen. Book review: SPSS in Practice. An Illustrated Guide. 2nd edn. Pasant K. Puri, Arnold Publishers, London, 2002. No. of pages: x + 175. ISBN 0-340-76112-1. Statistics in Medicine, 22(18):2964–2965, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lemmon:2005:DSS

[502] David R. Lemmon and J. L. (Joseph L.) Schafer. Developing statistical software in Fortran 95. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-28123-1, 0-387-23817-4 (paperback). xv + 323 pp. LCCN QA76.5

.L453 2005. URL http://www.loc.
gov/catdir/enhancements/fy0662/
2004061447-d.html;   http://www.
loc.gov/catdir/enhancements/fy0818/
2004061447-t.html.

Wade:2005:BRH

[503] Angie Wade. Book review: 1. A handbook of statistical analyses using SPSS. Sabine Landau and Brian S. Everitt, Chapman and Hall/CRC, Boca Raton, 2004. No. of pages: xii + 354. price: $44.95, £24.99. ISBN: 1-58488-369-3. Statistics in Medicine, 24(20): 3236–3237, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:BRHb

[504] Anonymous. Book review: Handbook of Univariate and Multivariate Data Analyses and Interpretation with SPSS, edited by Ho, R. Biometrics, 62(4): 1279–1280, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MarquesdeSa:2007:ASU

[505] Joaquim P. Marques de Sá. Applied Statistics Using SPSS, STATIS-TICA, MATLAB and R. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71972-5. xxiv + 506 + 196 pp. LCCN QA276.4 S2232 2007.

Anonymous:2009:BRRb

[506] Anonymous. Book review: R for SAS and SPSS Users, by Muenchen, R. A. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Privitera:2011:SBS

[507] Gregory J. Privitera. Statistics for the behavioral sciences. SAGE, Thousand Oaks, CA, USA, 2011. ISBN 1-4129-6931-X (hardcover). xxxix + 664 pp. LCCN HA29 .P755 2012.

Basto:2012:SRM

[508] Mário Basto and José Manuel Pereira.

An SPSS R-menu for ordinal factor analysis. Journal of Statistical Software, 46(4):??, January 2012. CODEN JS-SOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v46/i04.

Godfrey:2014:SRS

[509] A. Jonathan R. Godfrey and M. Theodor Loots. Software review: Statistical Software (R, SAS, SPSS, and Minitab) for Blind Students and Practitioners. Journal of Statistical Software, 58(SR-1):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v58/s01.

Ho:2014:HUM

[510] Robert Ho. Handbook of univariate and multivariate data analysis with IBM SPSS. Taylor and Francis, Boca Raton, FL, USA, second edition, 2014. ISBN 1-4398-9021-8 (hardcover). xxiii + 561 pp. LCCN QA279 .H6 2014.

Nooraee:2014:GLO

[511] Nazanin Nooraee, Geert Molenberghs, and Edwin R. van den Heuvel. GEE for longitudinal ordinal data: Comparing R-geepack, R-multgee, R-repolr, SAS-GENMOD, SPSS-GENLIN. Computational Statistics & Data Analysis, 77 (??):70–83, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000863.

Salkind:2014:SPW

[512] Neil J. Salkind. Statistics for people who (think they) hate statistics. SAGE Publications, Inc., Thousand Oaks, CA, USA, fifth edition, 2014. ISBN 1-4522-7771-0 (paperback), 1-4833-1288-7 (web PDF). xxvii + 483 pp. LCCN HA29 .S2365 2014.

Shanmugam:2015:BRSc

[513] Ramalingam Shanmugam. Book review: SPSS book is reviewed. [Building SPSS graphs to understand data, ISBN-13 978-1-4522-1648-3]. Journal of Statistical Computation and Simulation, 85(5):1070, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:EDM

[514] Jianjun Wang and Dallas E. Johnson. An examination of discrepancies in multiple imputation procedures between SAS^{T M} and SPSS^{T M}. The American Statistician, 73(1):80–88, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/
00031305.2018.1437078.

Zhao:2020:UAI

[515] Jie Zhao. User acceptance of information feed advertising: a hybrid method based on SEM and QCA. Future Internet, 12 (12):209, November 26, 2020. CODEN ???? ISSN 1999-5903. URL https://

www.mdpi.com/1999-5903/12/12/209.

Ortiz-Marcos:2021:CAI

[516] José Manuel Ortiz-Marcos. Cyberbullying analysis in intercultural educational environments using binary logistic regressions. Future Internet, 13(1):15, January 09, 2021. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.
com/1999-5903/13/1/15.

Hossain:2022:AUB

[517] Md. Moyazzem Hossain. Applied univariate, bivariate, and multivariate statistics: Understanding statistics for social and natural scientists, with applications in SPSS and R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 727–728, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2022:BRA

[518] Md. Moyazzem Hossain. Book review:

Applied univariate, bivariate, and multivariate statistics: Understanding statistics for social and natural scientists, with applications in SPSS and R, Daniel J. Denis, 2021 Hoboken, NJ, USA,, John Wiley & Sons, Inc. xxi + 551 pp., $104.00 (e-book), $130.00 (hardcover), ISBN: 978-1-119-58304-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 727–728, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

SPSS:1977:PAS

[519] Proceedings of the … Annual SPSS Users and Coordinators Conference. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1977.

SPSS:1978:PIS

[520] Proceedings of ISSUE ’78, the Second Annual SPSS Users and Coordinators Conference: Chicago, IL, October 8–11, 1978. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1978.

SPSS:1981:PIF

[521] Proceedings of ISSUE ’81: the Fifth Annual SPSS Users and Coordinators Conference, San Francisco, California, October 25–28, 1981. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1981.

SPSS:1982:PIS

[522] Proceedings of ISSUE ’82: the Sixth Annual SPSS Users and Coordinators Conference, New Orleans, Louisiana, October 10-14, 1982. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1982.

SPSS:1984:PIC

[523] Proceedings of the 1984 ISSUE Conference: Chicago, Illinois, 15–18 July 1984. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1984.

Allen:1986:CSS

[524] David M. Allen, editor. Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17– 19, 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4

.S95 1985. URL http://catalog.
hathitrust.org/api/volumes/oclc/
13455449.html;   http://www.dtic.
mil/dtic/tr/fulltext/u2/a166019.
pdf;   http://zbmath.org/?q=an:
0647.62008.